### v1

# FaithUn: Toward Faithful Forgetting in Language Models by Investigating the Interconnectedness of Knowledge

Abstract
Various studies have attempted to remove sensitive or private knowledge from a language model to prevent its unauthorized exposure. However, prior studies have overlooked the complex and interconnected nature of knowledge, where related knowledge must be carefully examined. Specifically, they have failed to evaluate whether an unlearning method faithfully erases interconnected knowledge that should be removed, retaining knowledge that appears relevant but exists in a completely different context. To resolve this problem, we first define a new concept called superficial unlearning, which refers to the phenomenon where an unlearning method either fails to erase the interconnected knowledge it should remove or unintentionally erases irrelevant knowledge. Based on the definition, we introduce a new benchmark, FaithUn, to analyze and evaluate the faithfulness of unlearning in real-world knowledge QA settings. Furthermore, we propose a novel unlearning method, KLUE, which updates only knowledge-related neurons to achieve faithful unlearning. KLUE identifies knowledge neurons using an explainability method and updates only those neurons using selected unforgotten samples. Experimental results demonstrate that widely-used unlearning methods fail to ensure faithful unlearning, while our method shows significant effectiveness in real-world QA unlearning.
FaithUn: Toward Faithful Forgetting in Language Models by Investigating the Interconnectedness of Knowledge
Nakyeong Yang1, Minsung Kim1, Seunghyun Yoon2, Joongbo Shin3 and Kyomin Jung1 1Seoul National University, 2Adobe Research, 3LG AI Research {yny0506, kms0805, kjung}@snu.ac.kr syoon@adobe.com, jb.shin@lgresearch.ai
1 Introduction
Large language models (LLMs) are trained on a vast corpus of text, enabling them to achieve outstanding performance across various tasks (Radford et al., 2019; Chowdhery et al., 2023; Gemma et al., 2024). However, LLMs may present privacy risks, as sensitive or private information could unintentionally be included in the large text corpus used for training. Therefore, prior studies have investigated unlearning undesirable knowledge in language models. To assess unlearning, most studies have examined whether a model successfully forgets the targeted knowledge while retaining irrelevant knowledge (Shi et al., 2024; Li et al., 2024a; Maini et al., 2024; Jin et al., 2024).
Figure 1: Faithful Unlearning. FaithUn proposes three types of datasets to evaluate the faithfulness of unlearning methods (i.e., Paraphrased, Multi-hop, and Same-answer datasets). Each target knowledge to be unlearned is mapped with questions corresponding to these three dataset types for evaluation.
However, they are limited since they have overlooked the complex and interconnected nature of knowledge, which requires careful investigation of related knowledge. Specifically, these studies have examined only the independent knowledge and failed to evaluate whether an unlearning method effectively erases interconnected knowledge that should be removed, while retaining knowledge that appears relevant but exists in a completely different context. Figure 1 presents an example of faithful unlearning in the real-world knowledge setting. Unlearning methods should also remove paraphrased and multi-hop questions, as they involve knowledge interconnected with the target question being unlearned. Conversely, unlearning methods should retain knowledge of other questions with the same answer as the target, if they actually contain different knowledge despite appearing relevant.
To address this gap, we first define superficial unlearning, which refers to the phenomenon where an unlearning method either fails to erase the interconnected knowledge it should remove or unintentionally erases irrelevant knowledge. Based on the definition, we introduce FaithUn (A Faithful Unlearning Evaluation Benchmark for Real-world Knowledge Question Answering), a new benchmark to examine three types of unlearning challenges: generalization, multi-hop unlearning, and shortcut unlearning to investigate superficial unlearning. Generalization (Anil et al., 2022; Yang et al., 2024a; Albalak et al., 2024), the multi-hop reasoning (Zhong et al., 2023; Li et al., 2024b; Yang et al., 2024b), and shortcut learning (Du et al., 2023; Tang et al., 2023; Zhou et al., 2023) are crucial challenges in machine learning research. Since the unlearning process typically relies on fewer data instances than general training, these challenges can be further amplified. Therefore, we construct three types of new datasets—paraphrased, multi-hop, and same-answer datasets—to examine superficial unlearning. These datasets address generalization, multi-hop knowledge unlearning, and shortcut unlearning, respectively. We demonstrate that existing unlearning methods do not ensure faithful unlearning, which raises new research questions for knowledge unlearning.
Furthermore, we propose a novel method, KLUE, which stands for Knowledge-Localized UnlEarning, to achieve faithful unlearning by precisely identifying and updating neurons related to the target knowledge. Specifically, we use attribution (Yang et al., 2023), an explainability method, to determine which neurons should be updated by quantifying how much information each neuron contributes to predicting the answer to a given question. However, the quantified score may include superficial knowledge that simply affects the target output’s probability without considering contextual meaning. Therefore, we propose a robust knowledge regularization method that accurately quantifies each neuron’s knowledge score, mitigating the superficial contribution of neurons. After identifying knowledge neurons, our method selectively unlearns the target knowledge while preserving other knowledge by updating only knowledge-related neurons with selected unforgotten samples. In our experiments, our method significantly outperforms the baselines in the FaithUn setting, demonstrating that knowledge-localized unlearning effectively achieves faithful unlearning.
2 Unlearning in Large Language Models
Machine unlearning has been used as a solution to address privacy and copyright issues in the text generation process of language models. Notable examples include gradient ascent-based methods (Jang et al., 2023; Yao et al., 2023; Barbulescu and Triantafillou, 2024), preference optimization approaches (Rafailov et al., 2024; Zhang et al., 2024; Jin et al., 2024), and representation learning techniques (Li et al., 2024a; Yao et al., 2024).
However, the effectiveness of these methods has not been clearly demonstrated, prompting prior studies to introduce benchmarks in the field of unlearning to assess them. Eldan and Russinovich (2023); Shi et al. (2024); Tian et al. (2024) have aimed to unlearn the knowledge of copyrighted texts (e.g., BBC News and Harry Potter book) in a language model. Li et al. (2024a) have introduced a benchmark dealing with hazardous knowledge in various professional domains (e.g., biosecurity and cybersecurity). Maini et al. (2024); Jin et al. (2024) have proposed benchmarks for unlearning various entities. Specifically, Maini et al. (2024) have created synthetic entity profiles and removed their knowledge from a language model. Jin et al. (2024) have tried to unlearn the knowledge about real-world entities and evaluated the knowledge memorization in various forms of assessment (e.g., cloze test and question answering). However, existing studies remain limited as they have only examined independent knowledge and overlooked the intricate nature of world knowledge. World knowledge is highly complex and interconnected, which means that unlearning the target knowledge requires examining related knowledge carefully. Our research focuses on this aspect, examining and facilitating faithful unlearning.
3 The FaithUn Benchmark
3.1 Problem Definition
The FaithUn task evaluates unlearning algorithms under real-world knowledge QA settings. Formally, given a language model $P_{\theta}(y|x)=\prod_{t=1}^{T}P_{\theta}(y_{t}|x,y_{1},...,y_{t-1})$ with parameters $\theta$, an unlearning algorithm $f$ updates $\theta$ to $\theta^{\prime}$, erasing the target knowledge from $P_{\theta}$. FaithUn includes various question-answer pairs $(q,a)\in\mathcal{C}$, where $\mathcal{C}$ is a question-answer pair set. Our task provides forget set $\mathcal{C}_{f}$, which contains target question-answer pairs to be forgotten, where $\mathcal{C}_{f}\subset\mathcal{C}$. FaithUn also provides retain set $\mathcal{C}_{r}\subset\mathcal{C}\backslash\mathcal{C}_{f}$ and test set $\mathcal{C}_{t}\subset\mathcal{C}\backslash(\mathcal{C}_{f}\cup\mathcal{C}_{r})$. $\mathcal{C}_{r}$ is used in the unlearning process as training samples to maintain the original knowledge of $P_{\theta}$, and $\mathcal{C}_{t}$ is used as unseen data to evaluate an unlearned model $P_{\theta^{\prime}}$ to reveal whether the unlearned model maintains the original knowledge. Furthermore, FaithUn provides other new types of test sets (i.e., paraphrased, multi-hop, and same-answer sets) to assess the faithfulness of unlearning methods. Before introducing the other datasets, we first define key aspects of our benchmark.
MUSE KnowUnDo WMDP TOFU RWKU FaithUn (Ours)
Knowledge Source News & Book Copyrighted books Hazardous knowledge Fictitious Author Real-world Entity Real-world Entity
# Unlearning Entities N/A N/A N/A 200 200 200
# Forget Probes 889 987 4,157 4,000 13,131 8,377
Knowledge Exists in LLMs X X O X O O
Generalization Test X X X X O O
Multi-hop Unlearning Test X X X X X O
Shortcut Unlearning Test X X X X X O
Table 1: Dataset Comparison. FaithUn aims to examine three types of unlearning challenges: generalization, multi-hop knowledge unlearning, and shortcut unlearning to investigate superficial unlearning. FaithUn can be used flexibly as it evaluates the removal of pre-existing knowledge about famous figures within LLMs.
World Knowledge Graph.
A world knowledge graph $\mathcal{K}$ is a directed multi-graph where nodes are entities and edges are labeled with relations, i.e., elements of two sets $\mathcal{E}$ and $\mathcal{R}$, respectively. We define $\mathcal{K}$ as a collection of triples $(s,r,o)\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}$, where $s$, $r$, $o$ denote the subject, relation, and object, respectively (Ruffinelli et al., 2020; Loconte et al., 2024). We assume that a world knowledge question is mapped to triples of $\mathcal{K}$; thus, we also define a knowledge mapping function, $\tau:\mathcal{Q}\rightarrow\mathcal{P}(\mathcal{K})$, where $\mathcal{Q}$ is a set of questions and $\mathcal{P}(\mathcal{K})$ represents the power set of $\mathcal{K}$. For example, the knowledge of a multi-hop question, $q_{i}$ = "Which continent is Tom Cruise’s country in?", can be denoted as a set of triples like $\kappa_{i}=\tau(q_{i})$ = {("Tom Cruise", "country", "U.S.A"), ("U.S.A", "continent", "North America")}.
To quantify memorization after unlearning, we define knowledge memorization of a language model following the general QA task, as follows:
Knowledge Memorization.
Let $P_{\theta}$ be a language model, and let $a$ be the correct answer to the question $q$. Then, knowledge memorization $\mathcal{M}_{\theta}:\mathcal{Q}\times\mathcal{A}\rightarrow\{0,1\}$ is defined as
[…]
where $\iota$ is an input prompt template for the language model $P_{\theta}$, and $\mathcal{Q}$ and $\mathcal{A}$ are question and answer sets, respectively. From the definition, $\mathcal{M}_{\theta}(q,a)=1$ indicates that the language model retains the knowledge of $(q,a)$, while $\mathcal{M}_{\theta}(q,a)=0$ signifies that it does not.
Furthermore, we define Superficial Unlearning using Knowledge Memorization as follows:
Superficial Unlearning.
Let $g:\Theta\rightarrow\Theta$ be an unlearning algorithm, and $\tau$ represent the knowledge mapping. Assume there is a forget set $\mathcal{C}_{f}$, where $\mathcal{M}_{\theta}(q,a)=1$ holds for all $(q,a)\in\mathcal{C}_{f}$, and that $(q_{j},a_{j})\notin\mathcal{C}_{f}$ with $\mathcal{M}_{\theta}(q_{j},a_{j})=1$. Furthermore, suppose we unlearn the knowledge of $\mathcal{C}_{f}$ using $g$ from a language model $P_{\theta}$, and finally get an unlearned model $P_{\theta^{\prime}}$. Then, $g$ is called a superficial unlearning algorithm for $\mathcal{C}_{f}$ if
$((\kappa_{f}\cap\kappa_{j}\neq\text{Ø})\>\wedge\>\mathcal{M}_{\theta^{\prime}}(q_{j},a_{j})=1)$ (2)
$\>\>\>\>\>\>\>\>\ \vee((\kappa_{f}\cap\kappa_{j}=\text{Ø})\wedge\mathcal{M}_{\theta^{\prime}}(q_{j},a_{j})=0),$
where $\kappa_{f}=\bigcup_{(q,a)\in\mathcal{C}_{f}}{\tau(q)}$ and $\kappa_{j}=\tau(q_{j})$.
For example, suppose that an unlearning algorithm $g$ unlearns the knowledge of the question $q_{i}$ = "Which country is Tom Cruise from?", but it does not unlearn the multi-hop question $q_{j}$ = "Which continent is Tom Cruise’s country in?". Then, the knowledge of two questions can be denoted as a set of knowledge triples like $\kappa_{i}=\tau(q_{i})=$("Tom Cruise", "country", "U.S.A")} and $\kappa_{j}=\tau(q_{j})=$("Tom Cruise", "country", "U.S.A"), ("U.S.A", "continent", "North America")}. In this case, $g$ is called a superficial unlearning algorithm since $\kappa_{i}\cap\kappa_{j}\neq\text{Ø}$ and $\mathcal{M}_{\theta^{\prime}}(q_{j},a_{j})=1$ is true; thus, the equation 2 is satisfied.
Faithful Unlearning Benchmark.
Based on the definition of superficial unlearning, we construct three new types of datasets: paraphrased, multi-hop, and same-answer sets to investigate the phenomenon of superficial unlearning. The paraphrased set $\mathcal{C}_{p}^{i}$, multi-hop set $\mathcal{C}_{m}^{i}$, and same-answer set $\mathcal{C}_{s}^{i}$ is matched with each question-answer pair $(q_{i},a_{i})\in\mathcal{C}$. The paraphrased set includes the same context questions with varying textual forms to the matched target question; thus, we should unlearn $\mathcal{C}_{p}^{i}$ if a matched question-answer pair $(q_{i},a_{i})$ is included in the forget set $\mathcal{C}_{f}$. The multi-hop set includes multi-hop question-answer pairs interconnected with the target question. Therefore, we should also unlearn $\mathcal{C}_{m}^{i}$ if a mapped pair $(q_{i},a_{i})$ is included in $\mathcal{C}_{f}$. The same-answer set includes question-answer pairs where the questions are from different contexts but share the same answer as $a_{i}$; thus, we should maintain the knowledge of the same-answer set, although a matched pair $(q_{i},a_{i})$ is included in $\mathcal{C}_{f}$.
3.2 Data Collection and Construction
Data Source. We construct FaithUn using Wikidata (Vrandečić and Krötzsch, 2014), a knowledge base including knowledge triples $(s,r,o)$ matched with millions of entities. We first select 200 of the most famous people as the entity set $\mathcal{E}$ from The Most Famous People Rank  [^1] https://today.yougov.com, and manually select 19 common relations as the relation set $\mathcal{R}$. The selected relations are shown in Appendix A.3.1.
The Base QA dataset. We retrieve all the triples $(s,r,o)$ from Wikidata, where $s\in\mathcal{E}$ and $r\in\mathcal{R}$. Based on these triples, we use GPT-4o mini [^2] https://openai.com/ to generate natural language form questions using a prompt template shown in Figure 6. We use an object (i.e., $o$) of each triple as the answer for each generated question. The constructed Base QA dataset $\mathcal{C}$ is split into three types of datasets: forget set $\mathcal{C}_{f}$, retain set $\mathcal{C}_{r}$, and test set $\mathcal{C}_{t}$.
Evaluation of Unlearning Generalization. We also generate the Paraphrased QA dataset $\mathcal{C}_{p}$ to evaluate the generalization of an unlearning method. Each question-answer pair $(q,a)\in\mathcal{C}$ is matched with three paraphrased questions. The Paraphrased QA dataset is generated during the Base QA dataset construction process by making GPT-4o mini generate four different questions for each triple. We use the first question as a sample of the Base QA dataset and the others for the Paraphrased QA dataset. We have strictly checked whether there are the same texts in the generated four texts by examining the lexical overlap between texts.
Evaluation of Multi-hop Knowledge Unlearning. We construct the Multi-hop QA dataset $\mathcal{C}_{m}$ to investigate superficial unlearning. Each question-answer pair $(q,a)\in\mathcal{C}$ is matched with multi-hop questions. After constructing the triples of the Base QA dataset, we additionally retrieve a set of chain-of-triples $((s_{1},r_{1},o_{1}),(s_{2},r_{2},o_{2}))$ from Wikidata, where $s_{1}\in\mathcal{E}$ and $r_{1},r_{2}\in\mathcal{R}$ and $o_{1}=s_{2}$. For each chain-of-triples, we also generate natural language questions using GPT-4o mini with the prompt template shown in Figure 7. We strictly validate that $o_{1}$ and $o_{2}$ are not included in the questions using instructions.
Evaluation of Shortcut Unlearning. We further build the Same-answer QA dataset $\mathcal{C}_{s}$. Each question-answer pair $(q,a)\in\mathcal{C}$ is also matched with the same-answer but different-context questions. After constructing the triples of the Base QA dataset, we also retrieve other triples $(s^{\prime},r^{\prime},o)$ that share the same object (i.e., $o$) with each triple from the Base QA dataset, where $s^{\prime}\notin\mathcal{E}$. We also generate natural language form questions using GPT-4o mini with the same prompt template used in constructing the Base QA dataset.
3.3 Dataset Summary
Dataset Format. Each instance of the dataset is denoted as a tuple: $d=\langle\mathcal{C}^{i},\mathcal{C}_{p}^{i},\mathcal{C}_{m}^{i},\mathcal{C}_{s}^{i}\rangle$. The FaithUn dataset starts from a core factual triple $(s,r,o)$, which forms the knowledge of the Base QA dataset $\mathcal{C}^{i}$. There are also the Paraphrased QA dataset $\mathcal{C}_{p}^{i}$, based on the same triple, the Multi-hop QA dataset $\mathcal{C}_{m}^{i}$, which extends from the original triple $(s,r,o)$, and the Same-answer QA dataset $\mathcal{C}_{s}^{i}$, which shares the same answers as the Base QA dataset’s questions but has different contexts. Each of these datasets ($\mathcal{C}^{i}$, $\mathcal{C}_{p}^{i}$, $\mathcal{C}_{m}^{i}$, and $\mathcal{C}_{s}^{i}$) is composed of question-answer pairs $(q,a)$, and they also include false answer options to enable evaluation through Multiple-choice QA (MCQA). The details for the MCQA setting are described in Section 3.4. We also describe detailed examples in Table 10. In addition, we summarize the differences our benchmark addresses compared to existing benchmarks (Shi et al., 2024; Tian et al., 2024; Li et al., 2024a; Maini et al., 2024; Jin et al., 2024) in Table 1.
Dataset Statistics. After collecting triples of the Base QA dataset, we filter only triples including matched Multi-hop QA or Same-answer QA samples. Therefore, each QA instance in the Base QA dataset serves as a cluster for evaluating the faithfulness of unlearning methods. Consequently, we collect 664 QA pairs for the Base QA dataset. Each Base QA instance includes three paraphrased questions, for a total of 1,992 paraphrased QA instances in our dataset. FaithUn also include 1,714 instances for multi-hop QA datasets. Furthermore, our dataset includes 4,671 instances for the Same-answer QA dataset. The statistics of the constructed FaithUn datasets are shown in Table 4.
Dataset Quality. We adopt a ChatGPT variant to generate natural language questions, a commonly used and powerful approach, following existing studies (Shi et al., 2024; Jin et al., 2024; Maini et al., 2024). However, to further investigate the quality of the dataset, we conducted a human evaluation for the generated questions. Specifically, we recruited crowd workers fluent in English through the university’s online community and had them evaluate 800 generated natural language questions. The results revealed an error rate of 0%, confirming the reliability of our benchmark.
3.4 Evaluation Framework
To evaluate the faithfulness of unlearning methods, we first split the forget set $\mathcal{C}_{f}$, the retaining set $\mathcal{C}_{r}$, and the test set $\mathcal{C}_{t}$ from the entire Base QA dataset $\mathcal{C}$. Then, we train a language model to unlearn the forget set while maintaining knowledge of the retaining set. We further evaluate the unlearned model to the test set to assess knowledge retention for unseen data. In addition, we evaluate the unlearned model with the other constructed datasets (i.e., $\mathcal{C}_{p}$, $\mathcal{C}_{m}$, and $\mathcal{C}_{s}$) mapped to the forget and test sets to analyze the aspect of superficial unlearning.
Our unlearning framework consists of two types of input formats: (1) general QA format, and (2) multiple-choice QA (MCQA) format. We use the general QA format for unlearning and the MCQA format for evaluation. The general QA format inputs a question without an additional template, while the MCQA format uses a template that includes instructions and answer options. Suppose we aim to unlearn the knowledge of the question "Who is the mother of Barack Obama?", then we train a language model not to output the correct answer (i.e., "Stanley Ann Dunham") using only the question as an input. However, many users use a language model with various instruction templates, and an unlearned model should be evaluated in a stricter environment considering generalization. Furthermore, evaluating all possible answers to a question is one of the most challenging aspects of QA evaluation. Therefore, we utilize the MCQA form to evaluate an unlearned model. This makes it easier for LLMs to derive knowledge since they are given answer options; thus, it makes unlearning algorithms harder to apply. For this reason, we use the MCQA setting to evaluate unlearned models in more challenging and practical settings.
3.5 Evaluation Metrics
We propose various metrics to evaluate the basic unlearning performance and the superficial unlearning performance. We use exact match to calculate the score of all metrics. (1) Unlearning Accuracy (UA): We compute accuracy for the forget set $\mathcal{C}_{f}$ to evaluate the basic unlearning performance. (2) Extended Unlearning Accuracy (UA): We compute accuracy for the Paraphrased QA set $\mathcal{C}_{p}$ to evaluate the generalized unlearning performance. (3) Test Accuracy (TA): We compute accuracy for the test set $\mathcal{C}_{t}$ to evaluate whether knowledge of unseen instances is maintained after the unlearning process. (4) Same-answer Test Accuracy (SA): We compute accuracy for the Same-answer QA set $\mathcal{C}_{s}$ to analyze shortcut unlearning. An unlearning algorithm may only superficially degrade the probability of the answer regardless of context. (5) Multi-hop Test Accuracy (MA): We compute accuracy for $\mathcal{C}_{m}$ matched with each instance of $\mathcal{C}_{f}$ and $\mathcal{C}_{t}$ to evaluate whether the interconnected knowledge of instances is effectively unlearned. To derive the aggregated MA score, we first compute the individual accuracies, MAf for all $(q,a)\in\mathcal{C}_{m}$ mapped to $\mathcal{C}_{f}$ and MAt for all $(q,a)\in\mathcal{C}_{m}$ mapped to $\mathcal{C}_{t}$; then, we compute the aggregated score, MA, by averaging the scores, ($100-$MAf) and MAt. Although the number of samples in $\mathcal{C}_{t}$ is generally higher than in $\mathcal{C}_{f}$, we average the scores with equal weight, as we assume that unlearning samples in $\mathcal{C}_{f}$ is important due to significant privacy concerns. (6) Total Score (Score): We aggregate all the evaluation scores by averaging ($100-$UA), TA, SA, and MA, to present the overall performance.
4 Method: KLUE
Unlearning methods should erase only the knowledge associated with the target knowledge while preserving all other knowledge. In this section, we describe the method, KLUE, that identifies neurons contextually related to the target knowledge and updates only them during the unlearning process.
4.1 Quantifying Knowledge Relevance
4.1.1 Knowledge Quantification
We utilize an attribution method (Shrikumar et al., 2016) to extract the importance of neurons for specific world knowledge from language models. It is usually used to derive the importance of the input features (i.e., pixel, token) for performing a specific task, but Yang et al. (2023) expands the attribution formula to the importance of intermediate neurons in language models. Formally, suppose we have $P_{\theta}(y|x)=\prod_{t=1}^{T}P_{\theta}(y_{t}|x,y_{1},...,y_{t-1})$ that represents a language model. The contribution of an $i$-th neuron to the representation $h$ in a particular layer, in predicting an answer $a$ given a question $q$ using $P_{\theta}$, is defined as follows:
[…]
where $h^{l}$ means $l$-th token representation of $h$, and $\partial P_{\theta}(a|q)/\partial h^{l}_{i}$ is the gradient of $P_{\theta}(a|q)$ with respect to $h^{l}_{i}$. In this study, we use transformer variants for experiments; thus, activation scores and gradients of a specific layer are computed for each input token representation. Therefore, if an input text includes $L$ tokens, we have $L$ attribution scores for each neuron; thus, we aggregate attributions of tokens by using max aggregation to acquire a single neuron knowledge attribution $A^{(q,a)}_{i}(h)$.
4.1.2 Superficial Knowledge Regularization
Equation 3 computes the knowledge relevance of each neuron for a specific $(q,a)$ pair. However, this equation may include undesirable information that only serves to increase the likelihood of the answer $a$ regardless of the given context. To eliminate undesirable information from the computed attribution, we construct synthetic mismatched QA pairs $(q^{\prime},a)\in\mathcal{C}^{\prime}$, where the answers remain the same as the target answer $a$, while the questions are randomly sampled independently of the answer. Then, we compute the attribution score for each mismatched pair and average them. Since a question and an answer included in mismatched pairs are contextually irrelevant, the computed attribution corresponds to the degree that unconditionally increases the likelihood of the answer regardless of the context (superficial knowledge). Therefore, we can compute the final knowledge attribution, $\mathcal{I}$, containing only contextual knowledge by excluding the information of the mismatched attribution from the basic knowledge attribution as follows:
$S^{(q,a)}_{i}(h)=\sum_{(q^{\prime},a)\in\mathcal{C}^{{}^{\prime}}}{\tilde{A}^{(q^{\prime},a)}_{i}(h)}$, (4)
$\mathcal{I}^{(q,a)}_{i}(h)=A^{(q,a)}_{i}(h)-\alpha\times\frac{1}{N}\times S^{(q,a)}_{i}(h)$,
where $\mathcal{C}^{{}^{\prime}}$ is a set including mismatched question and answer pairs. $N$ is the number of mismatched samples, and $\alpha$ is a hyper-parameter to determine the magnitude of knowledge exclusion. $\tilde{A}$ means a negative value of $A$ is converted to the zero value. Since the negative values of the attribution score are negative contributions to a specific knowledge, it is reasonable to eliminate that unnecessary information. We use $\mathcal{C}^{f}$ and $\mathcal{C}^{r}$ as a pool to sample mismatched questions. Notice that this regularization enhances the quantification of contextual knowledge; thus, it can improve multi-hop reasoning and mitigate shortcut unlearning.
4.2 Unforgotten Sample-localized Unlearning
If we repeatedly unlearn samples that have already been sufficiently unlearned, it leads to overfitting in language models. Therefore, we select only the samples that are not completely forgotten in the unlearning process to preserve the generalization performance. Specifically, in each epoch’s unlearning process, we select and unlearn only questions that satisfy the knowledge memorization criteria (Described in Section 3.1).
4.3 Knowledge Neuron-localized Unlearning
After selecting unforgotten samples, we localize and update only the knowledge neurons corresponding to those selected samples in the language model. Specifically, we first compute gradients of parameters for the selected unforgotten samples. Then, we quantify the knowledge relevance of each neuron by using the equations 3 and 4, and sort neurons of the whole target layers by the knowledge relevance scores; then, we select the top-$n$ knowledge neurons. We finally mask gradients of the parameters for knowledge-irrelevant neurons to exclude them from the unlearning process. Suppose that a weight matrix $\mathbf{W}\in\mathbb{R}^{d\times k}$ is a linear matrix multiplication parameter of a language model, and the gradient computed for the parameter is $\nabla_{\mathbf{W}}\mathcal{L}=\partial\mathcal{L}/\partial\mathbf{W}$. Then, the gradient of $i$-th neuron (i.e., column) of the weight matrix after masking is denoted as $\nabla_{\mathbf{W}_{:,i}}\tilde{\mathcal{L}}=$ $\gamma$ $\odot$ $\nabla_{\mathbf{W}_{:,i}}\mathcal{L}$, where $\gamma\in\{\mathbf{0}_{d},\mathbf{1}_{d}\}$ and $\odot$ means the Hadamard product. We also can mask bias terms similar to the weight matrix. Notice that this method is model-agnostic since all neural networks consist of linear transformation layers.
5 Experiments
5.1 FaithUn Setups
Models. We adopt the instruction-tuned Gemma-2 (Gemma et al., 2024) models (2B & 9B) and the Llama-3.2 (Dubey et al., 2024) model (3B) to evaluate unlearning methods since they are among the latest open-source language models showing excellent performance.
Data. We sample 5% as the forget set and 10% as the retaining set from the Base QA dataset $\mathcal{C}$ since there are generally fewer samples to unlearn than to retain in real-world scenarios. More experiments on varying numbers of samples for the forget set are shown in Appendix B.5. We select 70% of $\mathcal{C}$ as the test set, guaranteeing it is completely separate from $\mathcal{C}_{f}$ and $\mathcal{C}_{r}$. For the MCQA evaluation (Section 3.4), we manually select the instruction and randomly sample two false answer options from possible answers for each relation $r$. The details of an example of the MCQA format and selecting false answer options are shown in Appendix B.1 and B.2, respectively. We also conduct experiments on various prompt templates, and the results are described in Appendix B.6.4.
Training Setups. When unlearning is applied to a language model, there is often a trade-off between unlearning knowledge (i.e., UA, UA, and MAf) and retaining the model’s overall knowledge (i.e., TA, SA, and MAt). Therefore, choosing the optimal model in the unlearning process is challenging since unlearning and retention are both important. For a fair comparison, we early stop the training procedure when UA0.33absent0.33$\leq$ 0.33≤ 0.33 is satisfied (random sampling from three options) to select the optimal model. More detailed experimental settings can be found in Appendix B.3.
5.2 Baselines
We adopt widely-used unlearning methods to evaluate the superficial unlearning: Gradient Ascent (GA), Gradient Ascent with a Retaining Loss (GAret), two Direct Preference Optimization variants (DPOmis and DPOrej), NPO (Zhang et al., 2024), and RMU (Li et al., 2024a). More details for the baselines are described in Appendix B.3. For KLUE, we select only 5% of neurons from Feed-forward networks for the knowledge neuron localization, and update them using general gradient ascent with retention loss. We also use $\alpha=10$ and $N=5$ for the Superficial Knowledge Regularization term. The experiments analyzing various hyper-parameters are shown in Section 5.5 and Appendix B.6.
Method UA ($\downarrow$) TA ($\uparrow$) SA ($\uparrow$) MA ($\uparrow$) Score ($\uparrow$)
Default 81.82 85.99 79.63 48.67 -
GA 36.02 48.92 37.19 48.34 49.61
GAret 34.01 77.58 66.51 53.21 65.82
DPOrej 41.75 68.96 63.58 49.67 60.11
DPOmis 37.03 65.01 51.69 52.89 58.14
NPO 38.72 60.84 52.77 49.50 56.10
RMU 46.12 79.02 67.74 53.05 63.42
KLUE 36.70 82.97 74.69 58.16 69.78
Table 2: Gemma-2 (2B) experimental results. We report the results of four metrics after unlearning the forget set (5%) in our settings. Bolded results indicate the best performance. We compute the accuracy over three trials and report the average accuracy.
  Case Method Questions for Forgetting Questions for Testing Label Logit Shift
1 GAret "Where was Michael Jordan born?" (Paraphrased QA) "What city is known as the birthplace of Michael Jordan?" Brooklyn 0.5699 $\rightarrow$ 0.3333
KLUE 0.5699 $\rightarrow$ 0.3333
2 GAret "What is the country of citizenship of Ellen DeGeneres?" (Multi-hop QA) "What currency is associated with the country of citizenship United States dollar 0.5756 $\rightarrow$ 0.5757
KLUE                         of Ellen DeGeneres?" 0.5756 $\rightarrow$ 0.2163
3 GAret "Where was Khloé Kardashian born?" (Same-answer QA) "Where was Jamie Grace born?" Los Angeles 0.5556 $\rightarrow$ 0.2641
KLUE 0.5556 $\rightarrow$ 0.5652
4 GAret "Who is the mother of Charles III of the United Kingdom?" (Same-answer QA) "Who is Prince Andrew, Duke of York’s mother?" Elizabeth II 0.4850 $\rightarrow$ 0.3333
KLUE 0.4850 $\rightarrow$ 0.4315
Table 3: Qualitative Analysis. GAret and KLUE are given the same questions to forget ($\mathcal{C}_{f}$) and test ($\mathcal{C}_{p}$, $\mathcal{C}_{m}$, and $\mathcal{C}_{s}$). Red texts indicate questions that should be forgotten, while blue texts should be retained. The "Label" and "Logit Shift" columns represent the golden labels for test questions and the logit changes corresponding to the labels, respectively. The underlined logit values indicate that the unlearning result is successful.
5.3 KLUE Mitigates Superficial Unlearning
We investigate superficial unlearning on all baselines with Gemma-2 (2B) in the FaithUn setting, as shown in Table 2. First, the default Gemma-2 model can correctly answer most questions, validating that FaithUn is well constructed. After the unlearning process, all baselines reach UA0.33absent0.33$\leq$ 0.33≤ 0.33, which validates that all methods can unlearn target knowledge. However, they fail to reliably remove implicit and interconnected knowledge, suggesting that their unlearning process is superficial. However, our method mitigates superficial unlearning and achieves faithful unlearning compared to other baselines, without significantly damaging the other knowledge to maintain (i.e., TA, SA, and MA). These results demonstrate that our method accurately identifies neurons relevant to contextual knowledge and successfully erases this knowledge. In addition, experiments on Gemma-2 (9B) and Llama-3.2 (3B) reveal that our method outperforms baselines, with results presented in Appendix B.4. We also conduct ablation studies for KLUE and demonstrate the validity of our proposed methods, as detailed in Appendix B.7.
5.4 KLUE is Robust to Unlearning Trade-off.
We demonstrate how the unlearning process affects other knowledge by plotting all scores from the Gemma-2 (2B) unlearning process against UA. As the UA score represents the progress of unlearning target knowledge (decreasing with unlearning), we can observe each method’s impact on other knowledge in Figure 2. All methods’ impact on the paraphrased questions (UA) shows a strong correlation with the UA score, suggesting that all methods pose robustness in dealing with different lexical forms (but hold the same meaning) of the questions. However, the baselines struggle to maintain other knowledge (TA and SA) and to forget interconnected knowledge (MA). In contrast, KLUE demonstrates robust unlearning performance by effectively forgetting interconnected knowledge and preserving other knowledge.
Figure 2: The relationship between UA and other metrics. The X-axis shows UA in descending order, and the Y-axis shows the accuracy of other metrics.
5.5 The Impact of Neuron Localization
We adopt varying ratios of neuron selection $p\in\{0.01,0.05,0.1\}$ to investigate the effect of the knowledge neuron on Gemma-2 (2B). Also, we conduct experiments for the random neuron selection (i.e., $p\in\{0.01,0.05\}$). As a result, we reveal that a neuron ratio of 0.05 or 0.1 contributes to achieving faithful unlearning, showing that random neuron selection more significantly triggers superficial unlearning.
Figure 3: The ratio of neuron localization. We plot the accuracy of each metric for varying ratios of neuron localization.
5.6 Qualitative Analysis
We conduct a qualitative analysis for KLUE and GAret on Gemma-2 (2B). Both KLUE and GAret successfully unlearn the paraphrased question ($\mathcal{C}_{p}$), degrading label logits to 0.33. However, GAret has difficulty in unlearning the multi-hop question ($\mathcal{C}_{m}$), while mistakenly unlearns the same-answer questions ($\mathcal{C}_{s}$). On the other hand, KLUE faithfully unlearns them, mitigating superficial unlearning.
6 Conclusion
Our research identifies the limitation of existing unlearning benchmarks, which have not explored the interconnectedness of knowledge. To overcome this issue, we define superficial unlearning and propose a new benchmark, FaithUn, for evaluating generalization, multi-hop knowledge unlearning, and shortcut unlearning. Using this benchmark, we empirically demonstrate that existing unlearning methods are vulnerable to superficial unlearning. Furthermore, we propose a novel knowledge-localized unlearning method, KLUE, and demonstrate that it outperforms existing unlearning methods, effectively mitigating superficial unlearning. Our paper first illuminates the phenomenon of superficial unlearning and raises a new research question for a deeper analysis of the unlearning field.
Limitations
FaithUn is constructed based on Wikidata and is designed to investigate the unlearning of knowledge about famous people for application in various language models. Although knowledge is more interconnected for well-known individuals, our benchmark does not examine a broader range of people. Additionally, our study focuses solely on erasing the target label, leaving the issue of hallucinations in the unlearning process as future work, in line with prior studies.
Ethical Considerations
Our benchmark includes the private information of famous people, retrieved from Wikidata. Although the information of famous people is prevalent on the World Wide Web, the misuse of these data may raise ethical concerns regarding privacy.
Appendix A FaithUn Details
A.1 Detailed Dataset Comparison
In this section, we propose detailed comparisons with existing datasets to show the novelty of our benchmark clearly. Our benchmark aims to unlearn the knowledge of famous real-world entities, which can be prevalent in various language models, to consider the most practical situation of knowledge unlearning. Furthermore, our benchmark deals with the complex and interconnected nature of world knowledge; thus, we introduce three types of unlearning evaluation aspects (Generalization, Multi-hop knowledge unlearning, and Shortcut unlearning) for more deep analysis of real-world knowledge unlearning.
In summary, MUSE, KnowUnDo, and TOFU require fine-tuning to inject knowledge before unlearning, which may reduce their practicality. Additionally, only RWKU and our benchmark address real-world entities as targets for unlearning. Furthermore, most existing benchmarks, except for RWKU and our benchmark, have not considered related knowledge. However, RWKU has not explored the interconnections between knowledge and shortcut unlearning problems, which have become increasingly significant in unlearning due to its reliance on a limited number of training instances. For example, RWKU includes a target text for unlearning: ’Please forget Stephen King, who is an American author, renowned as the "King of Horror"’. It also contains a related knowledge question: "’Who plays the character Jack Torrance in the film The Shining?’". While the two questions are somewhat related, they represent independent pieces of knowledge, as they are not interconnected like multi-hop questions. In conclusion, the main contribution of our benchmark lies in evaluating whether unlearning methods perform faithful unlearning while considering knowledge interconnectedness within the real-world entity unlearning setting.
A.2 Dataset Format
Our FaithUn benchmark includes four types of datasets: the Base QA dataset ($\mathcal{C}$), the Paraphrased QA dataset ($\mathcal{C}_{p}$), the Multi-hop QA dataset ($\mathcal{C}_{m}$), and the Same-answer QA dataset ($\mathcal{C}_{s}$). Each instance in the Base QA dataset is matched with instances in other datasets (i.e., Paraphrased QA, Multi-hop QA, and Same-answer QA) to examine the impact of unlearning on these datasets. Dataset statistics for the FaithUn benchmark are shown in Table 4. Examples in the FaithUn benchmark are shown in Table 10.
Type Usage # instances Avg # in each cluster
Base QA train & test 664 1
Paraphrased QA test 1,992 3
Multi-hop QA test 1,714 2.68
Same-answer QA test 4,671 7.03
Table 4: Dataset statistics. FaithUn includes questions to be forgotten, collectively referred to as the Base QA dataset. Each question in this dataset forms a cluster, and questions from other datasets (i.e., Paraphrased QA, Multi-hop QA, and Same-answer QA) are matched with those in the Base QA dataset, thereby being assigned to the corresponding cluster.
A.3 Details in Dataset Construction
A.3.1 Selected Entities and Relations.
We select 200 famous human entities and 19 relations appropriate for constructing knowledge triples from Wikidata. Specifically, we manually select mother, country, religion, founded by, highest point, country of citizenship, place of birth, position played on team / speciality, headquarters location, country of origin, native language, field of work, father, occupation, sport, capital, currency, location, continent as relations, which are widely-used relations to describe knowledge of human entities or other entities related to human (e.g., United States of America).
A.3.2 Dataset Analysis.
The Number of Data Instances for Each Entity.
We investigate the number of data instances (cluster) for each entity, as shown in Figure 4. The X-axis of the figure corresponds to the entity index, which is sorted in descending order of popularity. From this figure, we can confirm that our dataset maintains a balanced distribution of entities, regardless of popularity. The average number of data instances of each entity is 3.32, and the standard deviation is 1.25.
Figure 4: The number of data instances per entity. The X-axis of the figure corresponds to the entity index, which is sorted in descending order of popularity. The Y-axis means the number of questions to be unlearned for each entity.
The Frequency of Each Relation.
we plot the number of each relation on the Base QA, the Multi-hop QA, and the Same-answer QA datasets, as shown in Figure 5. The Multi-hop QA dataset contains diverse relations, allowing for a broader evaluation of superficial unlearning. In contrast, the Same-answer QA dataset has a distribution of relation similar to the Base QA dataset, making unlearning more challenging. When evaluating shortcut unlearning on datasets with standardized relations, we can more effectively identify issues that lower the likelihood of predicting the given answer, regardless of context.
Figure 5: Relation frequency for each dataset. the Base QA dataset (left), the Multi-hop QA dataset (middle), and the Same-answer QA dataset (right).
A.3.3 Question Generation Prompt Templates
We utilize GPT-4o mini to generate questions from constructed Wikidata triples, similar to (Zhong et al., 2023; Mallen et al., 2022). An example of generating single-hop questions (the base QA, paraphrased QA, and same-answer QA datasets) is shown in Figure 6. Multi-hop questions are generated similarly to single-hop questions, shown in Figure 7.
Figure 6: Templates for generating single-hop questions using triples retrieved from Wikidata.
Figure 7: Templates for generating multi-hop questions using triples retrieved from Wikidata.
Appendix B Experimental Setup
B.1 MCQA Prompt Templates
The FaithUn framework evaluates unlearned models by using an MCQA format. The MCQA format consists of three parts: an instruction, a question, and options. After sampling false options for each question, we randomly shuffle the options to mitigate position bias (Pezeshkpour and Hruschka, 2024; Zheng et al., 2023), consistently maintaining the determined order during all the experiments for fair experiments. The utilized MCQA template is shown in Figure 8.
Figure 8: Templates for the multiple-choice question-answering (MCQA) prompting. We use this template to evaluate the knowledge of unlearned models accurately in a realistic usage scenario.
B.2 MCQA False Options Selection
To prevent the situation that the false options include a possible correct answer, we use GPT-4o  [^3] https://openai.com/index/hello-gpt-4o/ to cluster the entire answer options of each relation and we manually double-check the answer clusters are well constructed. After constructing answer clusters, we sample two incorrect options from the answer set, excluding those in the same cluster as the correct answer.
B.3 More Details for the Experiments
Training Setups.
We train and evaluate KLUE and other baselines on NVIDIA A100 GPU. For a fair comparison, we early stop the training procedure when UA0.33absent0.33$\leq$ 0.33≤ 0.33 is satisfied (random sampling from three answer options) to select the optimal model. Since a language model forgets all the knowledge when a learning rate is set too high, we have searched for the lowest learning rates, which can reach UA0.33absent0.33$\leq$ 0.33≤ 0.33 within the range $\lambda\in$ [1e-07, 3e-03]. We adopt batch size $\beta=4$ for all unlearning methods. We compute the final loss by weighted-summing the loss of forget samples and retaining samples. Specifically, we use $0.7$ and $1.0$ for the loss of forget samples and the retaining samples, respectively. We select $e=150$ as the maximum number of epochs in the training process.
Baselines.
(1) Gradient Ascent (GA): Unlike the gradient descent used during the pre-training phase, GA (Jang et al., 2023; Yao et al., 2023) maximize the negative log-likelihood loss on the forget set. This method helps shift the model away from its original predictions, aiding in the unlearning process. (2) Gradient Ascent with a Retaining Loss (GAret): GA tends to unlearn other unrelated knowledge since it just maximizes the negative log-likelihood loss on the forget set. Therefore, we add an auxiliary retention loss to maximize the log-likelihood of the retaining set, securing the retention of other irrelevant knowledge. (3) Direct Preference Optimization (DPO): We adopt preference optimization to unlearn a language model to generate another answer. DPO (Rafailov et al., 2024; Jin et al., 2024) utilizes positive and negative instances to train the model. Therefore, we select the correct answer as the negative instance and also define two types of DPO methods to determine positive ones: (1) DPOmis (DPO using a mismatched answer) and (2) DPOneg (DPO using a rejection answer). DPOmis utilizes a randomly sampled answer as the positive instance. On the other hand, DPOrej utilizes a rejection text “I can’t answer the question." as the positive instance. Two DPO methods both aim to increase the probability of the positive instance compared to the negative one for the forget set, and they switch the positive and negative instances for training the retaining set. We search for $\beta_{DPO}\in[0.1,0.5]$ to optimize models. (4) NPO: NPO is a modified version of DPO that exclusively retains negative examples without positive ones. NPO can also be explained as a straightforward modification of the GA loss. We implement NPO (Zhang et al., 2024) for extended experiments. We search for $\beta_{NPO}\in[0.1,0.5]$ to optimize models. (5) RMU: We implement RMU (Li et al., 2024a), the representation learning-based unlearning model. For RMU experiments, we search for $\alpha_{RMU}\in\{20,50,100,150,200,300\}$ and use hyper-parameters $c=20$ and $l=7$, following the implementation details on the original GitHub Page [^4] https://github.com/centerforaisafety/wmdp. (6) Knowledge-Localized Unlearning (KLUE): We select only 5% of neurons from Feed-forward networks for the knowledge neuron localization, and update them using general gradient ascent with retention loss. We also use $\alpha=10$ and $N=5$ for the Superficial Knowledge Regularization term. The experiments analyzing varying hyper-parameters are shown in Section 5.5, Appendix B.6.2, and Appendix B.6.3.
B.4 KLUE successfully Mitigates Superficial Unlearning on Various Language Models
We conduct experiments on Llama-3.2 (3B) and Gemma-2 (9B) to reveal that our method is model-agnostic and generalizable. Table 5 shows the experimental results for Llama-3.2 (3B) and Gemma-2 (9B), respectively. In the Llama-3.2 (3B) results, all baselines are significantly exposed to superficial unlearning, while our method successfully outperforms other baselines in most metrics. Likewise, in the Gemma-2 (9B) results, the enhancement of mitigating superficial unlearning is dramatic after applying knowledge localization. These results demonstrate that our method can be applied to various language models and successfully mitigates superficial unlearning.
Model Method UA ($\downarrow$) TA ($\uparrow$) SA ($\uparrow$) MA ($\uparrow$) Score ($\uparrow$)
Llama-3.2 (3B) Default 90.91 87.28 85.65 50.57 -
GA 35.35 54.52 39.19 52.45 52.70
GAret 48.14 68.24 57.71 53.94 57.94
DPOrej 46.80 69.68 55.86 54.02 58.19
DPOmis 36.02 64.87 43.21 51.56 55.91
KLUE 45.79 77.58 65.12 53.99 62.73
Gemma-2 (9B) Default 91.92 89.87 86.57 48.07 -
GA 29.29 40.52 30.56 50.46 48.06
GAret 45.45 83.84 68.52 50.72 64.40
DPOrej 41.41 75.32 59.72 47.02 60.16
DPOmis 36.36 63.15 43.06 55.45 56.32
KLUE 40.40 89.83 81.48 60.48 72.85
Table 5: Llama-3.2 (3B) and Gemma-2 (9B) experimental results. We report the results of four metrics after unlearning the forget set (5%) in our settings. Bolded results indicate the best performance.
B.5 KLUE is Robust to Various Forget Sample Sizes
We conduct experiments on Gemma-2 (2B) for the varying sizes (i.e., 1%, 5%, and 10%) of the forget set to analyze the effect of unlearning samples. The experimental results are shown in Table 2 (5%) and Table 6 (1% and 10%). Our experiments reveal that existing methods undergo more problems in unlearning when the number of forget samples increases. Increasing the number of samples to be forgotten is more challenging since it requires modifying a greater amount of knowledge. However, our proposed method consistently outperforms other baselines; thus, the performance gap between our method and the baselines widens as the number of forget samples increases.
Forget % Method UA ($\downarrow$) TA ($\uparrow$) SA ($\uparrow$) MA ($\uparrow$) Score ($\uparrow$)
1% Default 72.22 85.34 71.43 54.18 -
GA 44.44 77.80 57.14 49.43 59.98
GAret 34.33 85.78 59.52 58.38 67.33
DPOrej 44.44 72.84 54.76 51.79 58.73
KLUE 36.11 85.34 63.09 59.77 68.02
5% Default 81.82 85.99 79.63 48.67 -
GA 36.02 48.92 37.19 48.34 49.61
GAret 34.01 77.58 66.51 53.21 65.82
DPOrej 41.75 68.96 63.58 49.67 60.11
KLUE 36.70 82.97 74.69 58.16 69.78
10% Default 83.84 85.34 76.82 50.05 -
GA 38.38 28.02 31.13 50.41 42.79
GAret 40.40 62.50 65.12 54.21 60.35
DPOrej 34.85 45.26 42.38 51.29 51.02
KLUE 40.91 81.03 69.98 59.18 67.32
Table 6: Unlearning experiments for varying forget sample sizes. We report the unlearning results for the varying number of forget set (i.g., 1% and 10%). The results for 5% are also found in Table 2.
B.6 Hyper-parameter Experiments
B.6.1 Sequential vs. Batch Unlearning
We conduct experiments on Gemma-2 (2B) to show the performance variation for varying numbers of samples unlearned in each batch. We select 5% of neurons to unlearn. We adopt various batch size $\beta\in$1, 4, 8, 16, 32} for the experiments, shown in Figure 9. The experimental results reveal that KLUE is effective when using $\beta\in[4,16]$. Sequential unlearning restricts unlearning to specific knowledge for only a single data sample, which impacts overfitting in the unlearning process, resulting in good performance only on UA. In contrast, a large batch size makes it hard for a language model to unlearn the knowledge since it can not identify appropriate knowledge neurons from the attribution computed by large samples.
Figure 9: The batch size experiments.
B.6.2 Hyper-parameter ($\alpha$) Experiments
We conduct hyper-parameter experiments on Gemma-2 (2B) for $\alpha\in$0.5, 1.0, 10.0, 20.0}, which is used to determine the magnitude of the superficial knowledge regularization, shown in Figure 10. The experimental results show that low values of $\alpha$ damage the retention of the original knowledge (TA, SA), although they show better performance for unlearning interconnected knowledge of the forget set (UA). On the other hand, higher values of $\alpha$ contribute to preserving the retention of the original knowledge.
Figure 10: The hyper-param ($\alpha$) experiments.
B.6.3 Neuron Ratio ($p$) Experiments
We conduct experiments on various neuron ratios to investigate the KLUE method further for Gemma-2 (2B), as shown in Table 7. We reveal that even the larger ratios show comparable results, however, simply increasing the neuron ratio does not enhance the performance. The results also demonstrate that it is more important to exclude irrelevant neurons than to include relevant neurons during training to mitigate superficial unlearning.
Neurons ratio ($p$) UA TA SA MA Score
0.01 42.42 81.03 68.98 56.33 65.98
0.05 36.36 83.41 74.54 57.48 69.76
0.1 37.37 83.62 74.54 55.50 69.07
0.5 39.39 82.97 72.69 58.81 68.77
Table 7: The experiments on various neuron ratios.
B.6.4 The Various Prompt Templates Experiments
We conduct experiments on various prompt templates to investigate the unlearning abilities of the KLUE method further for Gemma-2 (2B), as shown in Table 7. Specifically, we newly select five templates: (1) "Pick the appropriate option for the question from the provided options. You should answer without further explanation.", (2) "Select the correct answer for the given question from the options. Write only the word without explanation.", (3) "Answer the given question by choosing the appropriate answer from the given options. Do not include any explanations.", (4) "Select the correct answer to the following question among the options. Only the exact word should be written, with no explanation.", and (5) "Select the proper answer to the question from among the given options. Write only the exact word without any additional explanation.". From the experiments, we reveal that the newly adopted prompts perform similarly to the original prompt. Their performance on the UA score is slightly higher than the original one since we early stopped the unlearning process based on the UA score evaluation for the original prompt.
prompt index UA UA TA SA MA Score
original 33.33 36.36 83.41 74.54 57.48 69.76
1 39.39 37.37 82.76 73.61 57.16 69.04
2 39.39 42.42 81.47 73.61 57.51 67.54
3 36.36 38.38 83.41 74.54 58.10 69.42
4 36.36 38.38 83.41 74.54 57.21 69.20
5 39.39 38.38 82.33 76.39 56.55 69.22
Table 8: The experiments on different prompts.
B.7 Ablation Studies
We perform ablation experiments on each KLUE method using Gemma-2 2B to better understand their relative importance, as shown in Table 9. Regularization means the strategy of using the auxiliary regularization term for quantifying the knowledge relevance of each neuron, mitigating superficial unlearning. Localization corresponds to the entire knowledge neuron localization strategy. Sample Selection is the strategy that selects unforgotten samples by evaluating the memorization of each sample. For the ablation study, we remove each of them and measure the accuracy. As a result, we reveal that three methods significantly affect the faithfulness of unlearning. Regularization and Localization are useful to enhance SA and MA, mitigating superficial unlearning related to interconnected knowledge. These results demonstrate that selecting proper knowledge neurons to be updated is useful for handling interconnected knowledge. In addition, we illuminate that Sample Selection significantly increases TA and SA, mitigating overfitting and shortcut unlearning issues.
Module UA ($\downarrow$) TA ($\uparrow$) SA ($\uparrow$) MA ($\uparrow$) Score ($\uparrow$)
Default 81.82 85.99 79.63 48.67 -
KLUE 36.70 82.97 74.69 58.16 69.78
(-) Regularization 40.40 79.74 67.59 51.24 64.54
(-) Localization 46.46 81.68 68.52 53.51 64.31
(-) Sample Selection 37.37 75.86 62.96 56.05 64.37
Table 9: Ablation studies
Type Notation Example
Example 1
Main triple $(s,r,o)$ (Hillary Clinton, father, Hugh E. Rodham)
Base QA $\mathcal{C}^{i}$ Who is the father of Hillary Clinton? $\rightarrow$ Hugh E. Rodham
False options: August Coppola, Earl Woods
Paraphrased QA $\mathcal{C}_{p}^{i}$ Who is Hillary Clinton’s dad? $\rightarrow$ Hugh E. Rodham
Who was Hillary Clinton’s father? $\rightarrow$ Hugh E. Rodham
What is the name of Hillary Clinton’s father? $\rightarrow$ Hugh E. Rodham
False options: August Coppola, Earl Woods
Multi-hop QA $\mathcal{C}_{m}^{i}$ What is the country of citizenship of Hillary Clinton’s father? $\rightarrow$ United States of America
False options: Spain, Vatican City
(Hillary Clinton, father, Hugh E. Rodham)
(Hugh E. Rodham, country of citizenship, United States of America)
What is the place of birth of Hillary Clinton’s father? $\rightarrow$ Scranton
False options: London, Pretoria
(Hillary Clinton, father, Hugh E. Rodham)
(Hugh E. Rodham, place of birth, Scranton)
Same-answer QA $\mathcal{C}_{s}^{i}$ Who is Anthony-Tony-Dean Rodham’s father? $\rightarrow$ Hugh E. Rodham
False options: Alfred Lennon, Hussein Onyango Obama
(Anthony-Tony-Dean Rodham, father, Hugh E. Rodham)
Example 2
Main triple $(s,r,o)$ (LeBron James, sport, basketball)
Base QA $\mathcal{C}^{i}$ What sport does LeBron James play? $\rightarrow$ basketball
False options: Auto racing, American football
Paraphrased QA $\mathcal{C}_{p}^{i}$ Which sport is associated with LeBron James? $\rightarrow$ basketball
In which sport is LeBron James a professional athlete? $\rightarrow$ basketball
What is the sport that LeBron James is known for? $\rightarrow$ basketball
False options: Auto racing, American football
Multi-hop QA $\mathcal{C}_{m}^{i}$ What is the country of origin of the sport that LeBron James plays? $\rightarrow$ United States of America
False options: Japan, Ryukyu Kingdom
(LeBron James, sport, basketball)
(basketball, country of origin, United States of America)
Same-answer QA $\mathcal{C}_{s}^{i}$ What sport does Kevin Durant play? $\rightarrow$ basketball
False options: Tennis, Boxing
(Kevin Durant, sport, basketball)
What sport is Wilt Chamberlain known for? $\rightarrow$ basketball
False options: Tennis, Auto racing
(Wilt Chamberlain, sport, basketball)
What sport is Larry Bird associated with? $\rightarrow$ basketball
False options: Association football, Aikido
(Larry Bird, sport, basketball)
Example 3
Main triple $(s,r,o)$ (Jackie Chan, place of birth, Victoria Peak)
Base QA $\mathcal{C}^{i}$ Where was Jackie Chan born? $\rightarrow$ Victoria Peak
False options: Jersey City, Louisiana
Paraphrased QA $\mathcal{C}_{p}^{i}$ What is the birthplace of Jackie Chan? $\rightarrow$ Victoria Peak
In which location was Jackie Chan born? $\rightarrow$ Victoria Peak
What place is known as the birth location of Jackie Chan? $\rightarrow$ Victoria Peak
False options: Jersey City, Louisiana
Multi-hop QA $\mathcal{C}_{m}^{i}$ What country is associated with the birthplace of Jackie Chan? $\rightarrow$ People’s Republic of China
False options: Australia, Mexico
(Jackie Chan, place of birth, Victoria Peak)
(Victoria Peak, country, People’s Republic of China)
Same-answer QA $\mathcal{C}_{s}^{i}$ Where was George Heath born? $\rightarrow$ Victoria Peak
False options: Neptune Township, Nuremberg
(George Heath, place of birth, Victoria Peak)
Where was Peter Hall born? $\rightarrow$ Victoria Peak
False options: Hawaii, Mission Hills
(Peter Hall, place of birth, Victoria Peak)
Table 10: Examples from the FaithUn dataset.

### v2

@@
 Various studies have attempted to remove sensitive or private knowledge from a language model to prevent its unauthorized exposure.
-However, prior studies have overlooked the complex and interconnected nature of knowledge, where related knowledge must be carefully examined.
-Specifically, they have failed to evaluate whether an unlearning method faithfully erases interconnected knowledge that should be removed, retaining knowledge that appears relevant but exists in a completely different context.
+However, prior studies have overlooked the inherent complexity and interconnectedness of knowledge, which requires careful examination.
 To resolve this problem, we first define a new concept called superficial unlearning, which refers to the phenomenon where an unlearning method either fails to erase the interconnected knowledge it should remove or unintentionally erases irrelevant knowledge.
-Based on the definition, we introduce a new benchmark, FaithUn, to analyze and evaluate the faithfulness of unlearning in real-world knowledge QA settings.
+Based on the definition, we introduce a novel benchmark, FaithUn, to analyze and evaluate the faithfulness of unlearning in real-world knowledge QA settings.
 Furthermore, we propose a novel unlearning method, KLUE, which updates only knowledge-related neurons to achieve faithful unlearning.
-KLUE identifies knowledge neurons using an explainability method and updates only those neurons using selected unforgotten samples.
-Experimental results demonstrate that widely-used unlearning methods fail to ensure faithful unlearning, while our method shows significant effectiveness in real-world QA unlearning.
+KLUE leverages a regularized explainability method to localize contextual knowledge neurons, updating only these neurons using carefully selected unforgotten samples.
+Experimental results demonstrate that existing unlearning methods fail to ensure faithful unlearning, while our method shows significant effectiveness in real-world QA unlearning.
 FaithUn: Toward Faithful Forgetting in Language Models by Investigating the Interconnectedness of Knowledge
-Nakyeong Yang1, Minsung Kim1, Seunghyun Yoon2, Joongbo Shin3 and Kyomin Jung1 1Seoul National University, 2Adobe Research, 3LG AI Research {yny0506, kms0805, kjung}@snu.ac.kr syoon@adobe.com, jb.shin@lgresearch.ai
+Nakyeong Yang1, Minsung Kim1, Seunghyun Yoon2, Joongbo Shin3, Kyomin Jung1 1Seoul National University, 2Adobe Research, 3LG AI Research {yny0506, kms0805, kjung}@snu.ac.kr syoon@adobe.com, jb.shin@lgresearch.ai
 1 Introduction
-Large language models (LLMs) are trained on a vast corpus of text, enabling them to achieve outstanding performance across various tasks (Radford et al., 2019; Chowdhery et al., 2023; Gemma et al., 2024).
-However, LLMs may present privacy risks, as sensitive or private information could unintentionally be included in the large text corpus used for training.
-Therefore, prior studies have investigated unlearning undesirable knowledge in language models.
-To assess unlearning, most studies have examined whether a model successfully forgets the targeted knowledge while retaining irrelevant knowledge (Shi et al., 2024; Li et al., 2024a; Maini et al., 2024; Jin et al., 2024).
-Figure 1: Faithful Unlearning.
-FaithUn proposes three types of datasets to evaluate the faithfulness of unlearning methods (i.e., Paraphrased, Multi-hop, and Same-answer datasets).
-Each target knowledge to be unlearned is mapped with questions corresponding to these three dataset types for evaluation.
-However, they are limited since they have overlooked the complex and interconnected nature of knowledge, which requires careful investigation of related knowledge.
-Specifically, these studies have examined only the independent knowledge and failed to evaluate whether an unlearning method effectively erases interconnected knowledge that should be removed, while retaining knowledge that appears relevant but exists in a completely different context.
-Figure 1 presents an example of faithful unlearning in the real-world knowledge setting.
-Unlearning methods should also remove paraphrased and multi-hop questions, as they involve knowledge interconnected with the target question being unlearned.
-Conversely, unlearning methods should retain knowledge of other questions with the same answer as the target, if they actually contain different knowledge despite appearing relevant.
-To address this gap, we first define superficial unlearning, which refers to the phenomenon where an unlearning method either fails to erase the interconnected knowledge it should remove or unintentionally erases irrelevant knowledge.
-Based on the definition, we introduce FaithUn (A Faithful Unlearning Evaluation Benchmark for Real-world Knowledge Question Answering), a new benchmark to examine three types of unlearning challenges: generalization, multi-hop unlearning, and shortcut unlearning to investigate superficial unlearning.
-Generalization (Anil et al., 2022; Yang et al., 2024a; Albalak et al., 2024), the multi-hop reasoning (Zhong et al., 2023; Li et al., 2024b; Yang et al., 2024b), and shortcut learning (Du et al., 2023; Tang et al., 2023; Zhou et al., 2023) are crucial challenges in machine learning research.
-Since the unlearning process typically relies on fewer data instances than general training, these challenges can be further amplified.
-Therefore, we construct three types of new datasets—paraphrased, multi-hop, and same-answer datasets—to examine superficial unlearning.
-These datasets address generalization, multi-hop knowledge unlearning, and shortcut unlearning, respectively.
+Large language models (LLMs) are trained on a vast corpus of text, enabling them to achieve outstanding performance across various tasks.
+However, LLMs may present privacy risks, as sensitive or private information may be inadvertently included in the text corpus used for training.
+Therefore, prior studies have examined unlearning undesirable knowledge in LLMs (Shi et al., 2024; Li et al., 2024; Maini et al., 2024; Jin et al., 2024; Lynch et al., 2024; Wu et al., 2024).
+Figure 1: Faithful Forgetting in LLMs.
+FaithUn proposes three datasets to evaluate unlearning methods (i.e., Paraphrased, Multi-hop, and Same-answer datasets).
+Each target knowledge to be unlearned is mapped with questions from these three datasets for evaluation.
+However, they are limited in that they have overlooked the complex and interconnected nature of knowledge, which necessitates a careful investigation of its internal dependencies.
+Figure 1 presents an example of faithful unlearning.
+Unlearning methods should also remove knowledge that is interconnected with the target questions to be unlearned—such as that found in paraphrased and multi-hop questions.
+Conversely, unlearning methods should retain knowledge that may appear relevant but is not directly connected to the target, such as questions that merely share the same answer.
+The unlearning process substantially relies on less data in training, as its goal is to remove only specific knowledge.
+Therefore, unlearned models tend to collapse into trivial solutions, unlike general training that utilizes large-scale data and enables broad generalization and multi-hop reasoning.
+   MUSE    KnowUnDo    WMDP    TOFU    RWKU    FaithUn (Ours)
+   Knowledge Source    News & Book    Copyrighted books    Hazardous knowledge    Fictitious Author    Real-world Entity    Real-world Entity
+   # Unlearning Entities    N/A    N/A    N/A    200    200    200
+   # Forget Probes    889    987    4,157    4,000    13,131    8,377
+   Knowledge Exists in LLMs
+   Generalization Test
+   Multi-hop Unlearning Test
+   Shortcut Unlearning Test
+Table 1: Dataset Comparison.
+FaithUn aims to examine three challenges: generalization, multi-hop knowledge unlearning, and shortcut unlearning to investigate superficial unlearning.
+FaithUn can be used flexibly to evaluate unlearning methods since it targets pre-existing knowledge of famous figures within LLMs.
+To address these problems, we first define superficial unlearning, which refers to the phenomenon where an unlearning method either fails to erase the interconnected knowledge it should remove or unintentionally erases irrelevant knowledge.
+Based on the definition, we introduce FaithUn (Faithful Unlearning Evaluation Benchmark), a new benchmark to investigate superficial unlearning.
+We construct three datasets—paraphrased, multi-hop, and same-answer—each addressing a key challenge: generalization, multi-hop knowledge unlearning, and shortcut unlearning, respectively.
 We demonstrate that existing unlearning methods do not ensure faithful unlearning, which raises new research questions for knowledge unlearning.
-Furthermore, we propose a novel method, KLUE, which stands for Knowledge-Localized UnlEarning, to achieve faithful unlearning by precisely identifying and updating neurons related to the target knowledge.
-Specifically, we use attribution (Yang et al., 2023), an explainability method, to determine which neurons should be updated by quantifying how much information each neuron contributes to predicting the answer to a given question.
+Furthermore, we propose a robust method, KLUE (Knowledge-Localized UnlEarning) to achieve faithful unlearning by precisely identifying and updating neurons related to the target knowledge.
+Specifically, we use the attribution method (Yang et al., 2023) to determine which neurons should be updated by quantifying how much each neuron contributes to predicting the answer to a given question.
 However, the quantified score may include superficial knowledge that simply affects the target output’s probability without considering contextual meaning.
-Therefore, we propose a robust knowledge regularization method that accurately quantifies each neuron’s knowledge score, mitigating the superficial contribution of neurons.
+Therefore, we propose a novel knowledge regularization method that accurately quantifies each neuron’s knowledge score, mitigating the trivial contribution of neurons.
 After identifying knowledge neurons, our method selectively unlearns the target knowledge while preserving other knowledge by updating only knowledge-related neurons with selected unforgotten samples.
-In our experiments, our method significantly outperforms the baselines in the FaithUn setting, demonstrating that knowledge-localized unlearning effectively achieves faithful unlearning.
-2 Unlearning in Large Language Models
-Machine unlearning has been used as a solution to address privacy and copyright issues in the text generation process of language models.
-Notable examples include gradient ascent-based methods (Jang et al., 2023; Yao et al., 2023; Barbulescu and Triantafillou, 2024), preference optimization approaches (Rafailov et al., 2024; Zhang et al., 2024; Jin et al., 2024), and representation learning techniques (Li et al., 2024a; Yao et al., 2024).
-However, the effectiveness of these methods has not been clearly demonstrated, prompting prior studies to introduce benchmarks in the field of unlearning to assess them.
-Eldan and Russinovich (2023); Shi et al.
-(2024); Tian et al.
-(2024) have aimed to unlearn the knowledge of copyrighted texts (e.g., BBC News and Harry Potter book) in a language model.
-Li et al.
-(2024a) have introduced a benchmark dealing with hazardous knowledge in various professional domains (e.g., biosecurity and cybersecurity).
-Maini et al.
-(2024); Jin et al.
-(2024) have proposed benchmarks for unlearning various entities.
-Specifically, Maini et al.
-(2024) have created synthetic entity profiles and removed their knowledge from a language model.
-Jin et al.
-(2024) have tried to unlearn the knowledge about real-world entities and evaluated the knowledge memorization in various forms of assessment (e.g., cloze test and question answering).
-However, existing studies remain limited as they have only examined independent knowledge and overlooked the intricate nature of world knowledge.
-World knowledge is highly complex and interconnected, which means that unlearning the target knowledge requires examining related knowledge carefully.
-Our research focuses on this aspect, examining and facilitating faithful unlearning.
+Our experiments reveal that existing methods fail to ensure faithful unlearning.
+However, KLUE significantly outperforms the baselines in the FaithUn setting, demonstrating that knowledge-localized unlearning effectively achieves faithful unlearning.
+In summary, this work makes the following contributions:
+  • We define superficial unlearning and introduce FaithUn, a new benchmark for evaluating whether unlearning methods can faithfully handle the interconnectedness of world knowledge.
+  • We reveal that existing methods fail to achieve faithful unlearning by showing a trivial solution, highlighting the need for further research.
+  • We propose KLUE, a knowledge-localized unlearning method that regularizes neuron attribution to identify and selectively update context-relevant neurons, achieving superior performance on FaithUn.
+2 Large Language Models Unlearning
+Machine unlearning has been used as a solution to address privacy and copyright issues in the text generation process of LLMs.
+Notable examples include gradient ascent-based methods (Jang et al., 2023; Yao et al., 2023; Barbulescu and Triantafillou, 2024), preference optimization approaches (Rafailov et al., 2024; Zhang et al., 2024; Jin et al., 2024), and representation learning techniques (Li et al., 2024; Yao et al., 2024).
+However, the effectiveness of these methods has not been clearly demonstrated, prompting the introduction of new benchmarks in the unlearning field to assess them.
+WHP, MUSE, and KnowUndo (Eldan and Russinovich, 2023; Shi et al., 2024; Tian et al., 2024) have aimed to unlearn the knowledge of copyrighted texts (e.g., News and Book).
+WMDP (Li et al., 2024) has introduced a benchmark for hazardous knowledge in professional domains (e.g., biosecurity).
+TOFU (Maini et al., 2024) has created synthetic profiles and removed their associated knowledge from LLMs.
+RWKU (Jin et al., 2024) has examined knowledge about real-world entities and evaluates memorization across various textual forms (e.g., cloze tests and QA) to assess generalization.
+While these studies have made valuable contributions, they fail to address the interconnected nature of knowledge.
+Even RWKU, despite its progress in surface-level generalization, fails to capture the deeper relational dependencies among pieces of knowledge.
+As a result, prior works have overlooked two critical issues—knowledge interconnections and shortcut unlearning—which are essential due to the inherently limited data available for unlearning.
+We summarize the differentiations of our benchmark compared to others in Table 1.
+Furthermore, we provide the detailed dataset comparisons in Appendix A.1.
 3 The FaithUn Benchmark
@@
 FaithUn includes various question-answer pairs $(q,a)\in\mathcal{C}$, where $\mathcal{C}$ is a question-answer pair set.
-Our task provides forget set $\mathcal{C}_{f}$, which contains target question-answer pairs to be forgotten, where $\mathcal{C}_{f}\subset\mathcal{C}$.
+Our task provides forget set $\mathcal{C}_{f}\subset\mathcal{C}$, which contains target question-answer pairs to be forgotten.
 FaithUn also provides retain set $\mathcal{C}_{r}\subset\mathcal{C}\backslash\mathcal{C}_{f}$ and test set $\mathcal{C}_{t}\subset\mathcal{C}\backslash(\mathcal{C}_{f}\cup\mathcal{C}_{r})$.
-$\mathcal{C}_{r}$ is used in the unlearning process as training samples to maintain the original knowledge of $P_{\theta}$, and $\mathcal{C}_{t}$ is used as unseen data to evaluate an unlearned model $P_{\theta^{\prime}}$ to reveal whether the unlearned model maintains the original knowledge.
+$\mathcal{C}_{r}$ is used in the unlearning process as training samples to maintain the original knowledge of $P_{\theta}$, and $\mathcal{C}_{t}$ is used as unseen data to evaluate an unlearned model $P_{\theta^{\prime}}$ to determine whether the unlearned model maintains the original knowledge.
 Furthermore, FaithUn provides other new types of test sets (i.e., paraphrased, multi-hop, and same-answer sets) to assess the faithfulness of unlearning methods.
 Before introducing the other datasets, we first define key aspects of our benchmark.
-MUSE KnowUnDo WMDP TOFU RWKU FaithUn (Ours)
-Knowledge Source News & Book Copyrighted books Hazardous knowledge Fictitious Author Real-world Entity Real-world Entity
-# Unlearning Entities N/A N/A N/A 200 200 200
-# Forget Probes 889 987 4,157 4,000 13,131 8,377
-Knowledge Exists in LLMs X X O X O O
-Generalization Test X X X X O O
-Multi-hop Unlearning Test X X X X X O
-Shortcut Unlearning Test X X X X X O
-Table 1: Dataset Comparison.
-FaithUn aims to examine three types of unlearning challenges: generalization, multi-hop knowledge unlearning, and shortcut unlearning to investigate superficial unlearning.
-FaithUn can be used flexibly as it evaluates the removal of pre-existing knowledge about famous figures within LLMs.
 World Knowledge Graph.
 A world knowledge graph $\mathcal{K}$ is a directed multi-graph where nodes are entities and edges are labeled with relations, i.e., elements of two sets $\mathcal{E}$ and $\mathcal{R}$, respectively.
-We define $\mathcal{K}$ as a collection of triples $(s,r,o)\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}$, where $s$, $r$, $o$ denote the subject, relation, and object, respectively (Ruffinelli et al., 2020; Loconte et al., 2024).
+We define $\mathcal{K}$ as a collection of triples $(s,r,o)\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}$, where $s$, $r$, $o$ denote the subject, relation, and object, respectively (Ruffinelli et al., 2020).
 We assume that a world knowledge question is mapped to triples of $\mathcal{K}$; thus, we also define a knowledge mapping function, $\tau:\mathcal{Q}\rightarrow\mathcal{P}(\mathcal{K})$, where $\mathcal{Q}$ is a set of questions and $\mathcal{P}(\mathcal{K})$ represents the power set of $\mathcal{K}$.
@@
 […]
-where $\iota$ is an input prompt template for the language model $P_{\theta}$, and $\mathcal{Q}$ and $\mathcal{A}$ are question and answer sets, respectively.
-From the definition, $\mathcal{M}_{\theta}(q,a)=1$ indicates that the language model retains the knowledge of $(q,a)$, while $\mathcal{M}_{\theta}(q,a)=0$ signifies that it does not.
+where $\iota$ is an input prompt template for the language model $P_{\theta}$, and $\mathcal{Q}$ and $\mathcal{A}$ are question and answer sets.
+Therefore, $\mathcal{M}_{\theta}(q,a)=1$ indicates that the model retains the knowledge of $(q,a)$, while $\mathcal{M}_{\theta}(q,a)=0$ signifies that it does not.
 Furthermore, we define Superficial Unlearning using Knowledge Memorization as follows:
@@
 $\>\>\>\>\>\>\>\>\ \vee((\kappa_{f}\cap\kappa_{j}=\text{Ø})\wedge\mathcal{M}_{\theta^{\prime}}(q_{j},a_{j})=0),$
-where $\kappa_{f}=\bigcup_{(q,a)\in\mathcal{C}_{f}}{\tau(q)}$ and $\kappa_{j}=\tau(q_{j})$.
-For example, suppose that an unlearning algorithm $g$ unlearns the knowledge of the question $q_{i}$ = "Which country is Tom Cruise from?", but it does not unlearn the multi-hop question $q_{j}$ = "Which continent is Tom Cruise’s country in?".
+where $\kappa_{f}=\bigcup_{(q,a)\in\mathcal{C}_{f}}{\hskip-1.42271pt\tau(q)}$ and $\kappa_{j}=\tau(q_{j})$.
+For example, suppose that an unlearning algorithm $g$ unlearns the question $q_{i}$ = "Which country is Tom Cruise from?", but it does not unlearn the multi-hop question $q_{j}$ = "Which continent is Tom Cruise’s country in?".
 Then, the knowledge of two questions can be denoted as a set of knowledge triples like $\kappa_{i}=\tau(q_{i})=$ {("Tom Cruise", "country", "U.S.A")} and $\kappa_{j}=\tau(q_{j})=$ {("Tom Cruise", "country", "U.S.A"), ("U.S.A", "continent", "North America")}.
-In this case, $g$ is called a superficial unlearning algorithm since $\kappa_{i}\cap\kappa_{j}\neq\text{Ø}$ and $\mathcal{M}_{\theta^{\prime}}(q_{j},a_{j})=1$ is true; thus, the equation 2 is satisfied.
+In this case, $g$ is called a superficial unlearning algorithm since $\kappa_{i}\cap\kappa_{j}\neq\text{Ø}$ and $\mathcal{M}_{\theta^{\prime}}(q_{j},a_{j})=1$ is true (11st condition).
+In another case, suppose $g$ unlearns only $\kappa_{i}$ = {("Tom Cruise", "country", "U.S.A")} but mistakenly also remove $\kappa_{j}$ = {("Andy Warhol", "country", "U.S.A")}.
+This satisfies superficial unlearning since $\kappa_{i}\cap\kappa_{j}=\text{Ø}$ and $\mathcal{M}_{\theta^{\prime}}(q_{j},a_{j})=0$ (22nd condition).
 Faithful Unlearning Benchmark.
@@
 The paraphrased set $\mathcal{C}_{p}^{i}$, multi-hop set $\mathcal{C}_{m}^{i}$, and same-answer set $\mathcal{C}_{s}^{i}$ is matched with each question-answer pair $(q_{i},a_{i})\in\mathcal{C}$.
-The paraphrased set includes the same context questions with varying textual forms to the matched target question; thus, we should unlearn $\mathcal{C}_{p}^{i}$ if a matched question-answer pair $(q_{i},a_{i})$ is included in the forget set $\mathcal{C}_{f}$.
+The paraphrased set includes the same context questions with varying textual forms to the matched target question; thus, we should unlearn them if a matched question-answer pair $(q_{i},a_{i})$ is included in the forget set.
 The multi-hop set includes multi-hop question-answer pairs interconnected with the target question.
-Therefore, we should also unlearn $\mathcal{C}_{m}^{i}$ if a mapped pair $(q_{i},a_{i})$ is included in $\mathcal{C}_{f}$.
-The same-answer set includes question-answer pairs where the questions are from different contexts but share the same answer as $a_{i}$; thus, we should maintain the knowledge of the same-answer set, although a matched pair $(q_{i},a_{i})$ is included in $\mathcal{C}_{f}$.
+Therefore, we should also unlearn them if a mapped pair $(q_{i},a_{i})$ is included in the forget set.
+The same-answer set includes question-answer pairs where the questions are from different contexts but share the same answer as $a_{i}$; thus, we should maintain the knowledge of the same-answer set, although a matched pair $(q_{i},a_{i})$ is included in the forget set.
 3.2 Data Collection and Construction
 Data Source.
-We construct FaithUn using Wikidata (Vrandečić and Krötzsch, 2014), a knowledge base including knowledge triples $(s,r,o)$ matched with millions of entities.
+We construct FaithUn using Wikidata (Vrandečić et al., 2014), a knowledge base including knowledge triples $(s,r,o)$ matched with millions of entities.
 We first select 200 of the most famous people as the entity set $\mathcal{E}$ from The Most Famous People Rank  [^1] https://today.yougov.com, and manually select 19 common relations as the relation set $\mathcal{R}$.
@@
 We retrieve all the triples $(s,r,o)$ from Wikidata, where $s\in\mathcal{E}$ and $r\in\mathcal{R}$.
-Based on these triples, we use GPT-4o mini [^2] https://openai.com/ to generate natural language form questions using a prompt template shown in Figure 6.
+Based on these triples, we use GPT-4o mini to generate natural language form questions using a prompt template shown in Figure 6.
+Note that converting triples into natural language questions is a simple task, and most LLMs are capable of performing it.
 We use an object (i.e., $o$) of each triple as the answer for each generated question.
-The constructed Base QA dataset $\mathcal{C}$ is split into three types of datasets: forget set $\mathcal{C}_{f}$, retain set $\mathcal{C}_{r}$, and test set $\mathcal{C}_{t}$.
-Evaluation of Unlearning Generalization.
-We also generate the Paraphrased QA dataset $\mathcal{C}_{p}$ to evaluate the generalization of an unlearning method.
+The constructed Base QA dataset is split into three types of datasets: forget set, retain set, and test set.
+Assessing Unlearning Generalization.
+We also generate the Paraphrased QA dataset to evaluate the generalization of an unlearning method.
 Each question-answer pair $(q,a)\in\mathcal{C}$ is matched with three paraphrased questions.
@@
 We use the first question as a sample of the Base QA dataset and the others for the Paraphrased QA dataset.
-We have strictly checked whether there are the same texts in the generated four texts by examining the lexical overlap between texts.
-Evaluation of Multi-hop Knowledge Unlearning.
-We construct the Multi-hop QA dataset $\mathcal{C}_{m}$ to investigate superficial unlearning.
+We have strictly checked whether there are the same texts in the four generated texts by examining the lexical overlap between texts.
+Assessing Multi-hop Unlearning.
+We construct the Multi-hop QA dataset to investigate superficial unlearning.
 Each question-answer pair $(q,a)\in\mathcal{C}$ is matched with multi-hop questions.
 After constructing the triples of the Base QA dataset, we additionally retrieve a set of chain-of-triples $((s_{1},r_{1},o_{1}),(s_{2},r_{2},o_{2}))$ from Wikidata, where $s_{1}\in\mathcal{E}$ and $r_{1},r_{2}\in\mathcal{R}$ and $o_{1}=s_{2}$.
-For each chain-of-triples, we also generate natural language questions using GPT-4o mini with the prompt template shown in Figure 7.
-We strictly validate that $o_{1}$ and $o_{2}$ are not included in the questions using instructions.
-Evaluation of Shortcut Unlearning.
-We further build the Same-answer QA dataset $\mathcal{C}_{s}$.
+For each chain-of-triples, we also generate questions using GPT-4o mini with the prompt template shown in Figure 7.
+We ensure that $o_{1}$ and $o_{2}$ are not included in the questions with an instruction, and validate this with the lexical overlaps.
+Assessing Shortcut Unlearning.
+We further build the Same-answer QA dataset.
 Each question-answer pair $(q,a)\in\mathcal{C}$ is also matched with the same-answer but different-context questions.
 After constructing the triples of the Base QA dataset, we also retrieve other triples $(s^{\prime},r^{\prime},o)$ that share the same object (i.e., $o$) with each triple from the Base QA dataset, where $s^{\prime}\notin\mathcal{E}$.
-We also generate natural language form questions using GPT-4o mini with the same prompt template used in constructing the Base QA dataset.
+We also generate questions using GPT-4o mini with the same prompt template used in constructing the Base QA dataset.
 3.3 Dataset Summary
-Dataset Format.
-Each instance of the dataset is denoted as a tuple: $d=\langle\mathcal{C}^{i},\mathcal{C}_{p}^{i},\mathcal{C}_{m}^{i},\mathcal{C}_{s}^{i}\rangle$.
-The FaithUn dataset starts from a core factual triple $(s,r,o)$, which forms the knowledge of the Base QA dataset $\mathcal{C}^{i}$.
-There are also the Paraphrased QA dataset $\mathcal{C}_{p}^{i}$, based on the same triple, the Multi-hop QA dataset $\mathcal{C}_{m}^{i}$, which extends from the original triple $(s,r,o)$, and the Same-answer QA dataset $\mathcal{C}_{s}^{i}$, which shares the same answers as the Base QA dataset’s questions but has different contexts.
-Each of these datasets ($\mathcal{C}^{i}$, $\mathcal{C}_{p}^{i}$, $\mathcal{C}_{m}^{i}$, and $\mathcal{C}_{s}^{i}$) is composed of question-answer pairs $(q,a)$, and they also include false answer options to enable evaluation through Multiple-choice QA (MCQA).
-The details for the MCQA setting are described in Section 3.4.
-We also describe detailed examples in Table 10.
-In addition, we summarize the differences our benchmark addresses compared to existing benchmarks (Shi et al., 2024; Tian et al., 2024; Li et al., 2024a; Maini et al., 2024; Jin et al., 2024) in Table 1.
 Dataset Statistics.
@@
 The statistics of the constructed FaithUn datasets are shown in Table 7.
+We also describe detailed examples in Table 12.
 Dataset Quality.
 We adopt a ChatGPT variant to generate natural language questions, a commonly used and powerful approach, following existing studies (Shi et al., 2024; Jin et al., 2024; Maini et al., 2024).
-However, to further investigate the quality of the dataset, we conducted a human evaluation for the generated questions.
+However, to further investigate the quality of the dataset, we conducted a human evaluation of the generated questions.
 Specifically, we recruited crowd workers fluent in English through the university’s online community and had them evaluate 800 generated natural language questions.
@@
 3.4 Evaluation Framework
-To evaluate the faithfulness of unlearning methods, we first split the forget set $\mathcal{C}_{f}$, the retaining set $\mathcal{C}_{r}$, and the test set $\mathcal{C}_{t}$ from the entire Base QA dataset $\mathcal{C}$.
-Then, we train a language model to unlearn the forget set while maintaining knowledge of the retaining set.
-We further evaluate the unlearned model to the test set to assess knowledge retention for unseen data.
-In addition, we evaluate the unlearned model with the other constructed datasets (i.e., $\mathcal{C}_{p}$, $\mathcal{C}_{m}$, and $\mathcal{C}_{s}$) mapped to the forget and test sets to analyze the aspect of superficial unlearning.
+To evaluate the faithfulness of unlearning methods, we first split the forget set, the retain set, and the test set from the entire Base QA dataset.
+Then, we train LLMs to unlearn the forget set while maintaining knowledge of the retain set.
+We further evaluate the unlearned model on the test set to assess the knowledge for unseen data.
+In addition, we evaluate it on other datasets—the Paraphrased, Multi-hop, and Same-answer QA datasets—mapped to the forget and test sets.
 Our unlearning framework consists of two types of input formats: (1) general QA format, and (2) multiple-choice QA (MCQA) format.
@@
 The general QA format inputs a question without an additional template, while the MCQA format uses a template that includes instructions and answer options.
-Suppose we aim to unlearn the knowledge of the question "Who is the mother of Barack Obama?", then we train a language model not to output the correct answer (i.e., "Stanley Ann Dunham") using only the question as an input.
-However, many users use a language model with various instruction templates, and an unlearned model should be evaluated in a stricter environment considering generalization.
-Furthermore, evaluating all possible answers to a question is one of the most challenging aspects of QA evaluation.
-Therefore, we utilize the MCQA form to evaluate an unlearned model.
+Suppose we aim to unlearn the knowledge of the question "Who is the mother of Barack Obama?", then we train an LLM not to output the correct answer (i.e., "Stanley Ann Dunham") using only the question as an input.
+However, many users use LLMs with various instruction templates, and an unlearned model should be evaluated in a stricter environment, considering generalization.
+Furthermore, assessing all possible answers to a question is one of the most challenging aspects of QA evaluation.
+Therefore, we utilize the MCQA form to assess an unlearned model.
 This makes it easier for LLMs to derive knowledge since they are given answer options; thus, it makes unlearning algorithms harder to apply.
-For this reason, we use the MCQA setting to evaluate unlearned models in more challenging and practical settings.
+For this reason, we use the MCQA form to assess unlearned models in more challenging and practical settings.
+The details for the MCQA setting are described in Appendix B.1 and B.2.
 3.5 Evaluation Metrics
-We propose various metrics to evaluate the basic unlearning performance and the superficial unlearning performance.
+We propose five metrics to evaluate the basic unlearning and the superficial unlearning performance.
 We use exact match to calculate the score of all metrics.
-(1) Unlearning Accuracy (UA): We compute accuracy for the forget set $\mathcal{C}_{f}$ to evaluate the basic unlearning performance.
-(2) Extended Unlearning Accuracy (UA): We compute accuracy for the Paraphrased QA set $\mathcal{C}_{p}$ to evaluate the generalized unlearning performance.
-(3) Test Accuracy (TA): We compute accuracy for the test set $\mathcal{C}_{t}$ to evaluate whether knowledge of unseen instances is maintained after the unlearning process.
-(4) Same-answer Test Accuracy (SA): We compute accuracy for the Same-answer QA set $\mathcal{C}_{s}$ to analyze shortcut unlearning.
-An unlearning algorithm may only superficially degrade the probability of the answer regardless of context.
-(5) Multi-hop Test Accuracy (MA): We compute accuracy for $\mathcal{C}_{m}$ matched with each instance of $\mathcal{C}_{f}$ and $\mathcal{C}_{t}$ to evaluate whether the interconnected knowledge of instances is effectively unlearned.
-To derive the aggregated MA score, we first compute the individual accuracies, MAf for all $(q,a)\in\mathcal{C}_{m}$ mapped to $\mathcal{C}_{f}$ and MAt for all $(q,a)\in\mathcal{C}_{m}$ mapped to $\mathcal{C}_{t}$; then, we compute the aggregated score, MA, by averaging the scores, ($100-$MAf) and MAt.
-Although the number of samples in $\mathcal{C}_{t}$ is generally higher than in $\mathcal{C}_{f}$, we average the scores with equal weight, as we assume that unlearning samples in $\mathcal{C}_{f}$ is important due to significant privacy concerns.
-(6) Total Score (Score): We aggregate all the evaluation scores by averaging ($100-$UA), TA, SA, and MA, to present the overall performance.
+(1) Unlearning Accuracy (UA): The accuracy for the forget set to evaluate the basic unlearning performance.
+(2) Extended Unlearning Accuracy (UA): The accuracy for the Paraphrased QA set to evaluate the generalized unlearning performance.
+(3) Test Accuracy (TA): The accuracy for the test set to evaluate whether knowledge of unseen data is maintained after the unlearning process.
+(4) Same-answer Test Accuracy (SA): The accuracy for the Same-answer QA set to analyze shortcut unlearning.
+An unlearning algorithm may only superficially degrade the probability of the answer regardless of context, as a trivial solution.
+(5) Multi-hop Test Accuracy (MA): The accuracy for the Multi-hop QA set matched with each instance of the forget set and test set to evaluate whether the interconnected knowledge of instances is effectively unlearned.
+We first derive individual accuracies for the multi-hop questions mapped to the forget set and test set, respectively.
+We denote MAf as the accuracy for the multi-hop questions mapped to the forget set, and MAt for the multi-hop questions mapped to the test set.
+Then, we compute the aggregated score, MA, by averaging the scores, (100100-MAf) and MAt.
+Although the number of samples in the test set is generally larger than in the forget set, we compute the average scores with equal weight, based on the assumption that unlearning the forget set is critical due to significant privacy concerns.
+(6) Total Score (Score): We average all the evaluation scores, (100100-UA), TA, SA, and MA, to present the overall performance.
 4 Method: KLUE
-Unlearning methods should erase only the knowledge associated with the target knowledge while preserving all other knowledge.
 In this section, we describe the method, KLUE, that identifies neurons contextually related to the target knowledge and updates only them during the unlearning process.
@@
 4.1.1 Knowledge Quantification
-We utilize an attribution method (Shrikumar et al., 2016) to extract the importance of neurons for specific world knowledge from language models.
-It is usually used to derive the importance of the input features (i.e., pixel, token) for performing a specific task, but Yang et al.
-(2023) expands the attribution formula to the importance of intermediate neurons in language models.
+We utilize an attribution method (Yang et al., 2023) to extract the importance of neurons for specific world knowledge from LLMs.
 Formally, suppose we have $P_{\theta}(y|x)=\prod_{t=1}^{T}P_{\theta}(y_{t}|x,y_{1},...,y_{t-1})$ that represents a language model.
@@
 […]
-[…]
 where $h^{l}$ means $l$-th token representation of $h$, and $\partial P_{\theta}(a|q)/\partial h^{l}_{i}$ is the gradient of $P_{\theta}(a|q)$ with respect to $h^{l}_{i}$.
-In this study, we use transformer variants for experiments; thus, activation scores and gradients of a specific layer are computed for each input token representation.
-Therefore, if an input text includes $L$ tokens, we have $L$ attribution scores for each neuron; thus, we aggregate attributions of tokens by using max aggregation to acquire a single neuron knowledge attribution $A^{(q,a)}_{i}(h)$.
+We use transformer variants for experiments; thus, activation scores and gradients of a specific layer are computed for each input token.
+Therefore, if an input text includes $L$ tokens, we have $L$ attribution scores for each neuron; thus, we aggregate attributions of tokens by using max pooling to acquire a single neuron attribution $A^{(q,a)}_{i}(h)$.
 4.1.2 Superficial Knowledge Regularization
-Equation 3 computes the knowledge relevance of each neuron for a specific $(q,a)$ pair.
-However, this equation may include undesirable information that only serves to increase the likelihood of the answer $a$ regardless of the given context.
-To eliminate undesirable information from the computed attribution, we construct synthetic mismatched QA pairs $(q^{\prime},a)\in\mathcal{C}^{\prime}$, where the answers remain the same as the target answer $a$, while the questions are randomly sampled independently of the answer.
-Then, we compute the attribution score for each mismatched pair and average them.
-Since a question and an answer included in mismatched pairs are contextually irrelevant, the computed attribution corresponds to the degree that unconditionally increases the likelihood of the answer regardless of the context (superficial knowledge).
-Therefore, we can compute the final knowledge attribution, $\mathcal{I}$, containing only contextual knowledge by excluding the information of the mismatched attribution from the basic knowledge attribution as follows:
-$S^{(q,a)}_{i}(h)=\sum_{(q^{\prime},a)\in\mathcal{C}^{{}^{\prime}}}{\tilde{A}^{(q^{\prime},a)}_{i}(h)}$, (4)
+We identify one of the primary reasons that unlearning methods fail to operate faithfully as their tendency to adopt a trivial solution, namely shortcut unlearning, which reduces the probability of the target answer without considering the context.
+If the attribution score is computed solely based on the original question-answer pair $(q,a)\in\mathcal{C}$ targeted for unlearning, there is a potential risk that the method may select neurons that simply increase the likelihood of the answer $a$, regardless of context.
+To address this, we introduce a novel method, superficial knowledge regularization, which effectively excludes neurons associated solely with the answer but grounded in irrelevant contexts.
+Specifically, we first construct synthetic mismatched QA pairs $(q^{\prime},a)\in\mathcal{C}^{\prime}$, where $q^{\prime}$ is randomly sampled without regard to $a$, while the answer remains the same as the target answer.
+Then, we compute and average the attribution scores across all mismatched pairs.
+Consequently, we derive the final knowledge attribution $\mathcal{I}$, which captures only contextual knowledge by subtracting the mismatched attribution from the basic attribution, as follows:
+$S^{(q,a)}_{i}(h)=\hskip-5.69046pt\sum_{(q^{\prime},a)\in\mathcal{C}^{{}^{\prime}}}{\hskip-7.11317pt\tilde{A}^{(q^{\prime},a)}_{i}(h)}$; (4)
 $\mathcal{I}^{(q,a)}_{i}(h)=A^{(q,a)}_{i}(h)-\alpha\times\frac{1}{N}\times S^{(q,a)}_{i}(h)$,
@@
 $\tilde{A}$ means a negative value of $A$ is converted to the zero value.
-Since the negative values of the attribution score are negative contributions to a specific knowledge, it is reasonable to eliminate that unnecessary information.
-We use $\mathcal{C}^{f}$ and $\mathcal{C}^{r}$ as a pool to sample mismatched questions.
-Notice that this regularization enhances the quantification of contextual knowledge; thus, it can improve multi-hop reasoning and mitigate shortcut unlearning.
-4.2 Unforgotten Sample-localized Unlearning
+Since the negative values of the attribution are negative contributions to specific knowledge, we eliminate that unnecessary information.
+We use the forget and retain sets as a pool to sample mismatched questions.
+This approach mitigates the risk of shortcut unlearning, thereby naturally aligning with the goals of contextual unlearning.
+Notice that alleviating unlearning behaviors that disregard context inherently aligns with the objective of contextual unlearning; thus, it can improve multi-hop reasoning and mitigate shortcut unlearning.
+4.2 Unforgotten Sample Unlearning
 If we repeatedly unlearn samples that have already been sufficiently unlearned, it leads to overfitting in language models.
-Therefore, we select only the samples that are not completely forgotten in the unlearning process to preserve the generalization performance.
-Specifically, in each epoch’s unlearning process, we select and unlearn only questions that satisfy the knowledge memorization criteria (Described in Section 3.1).
-4.3 Knowledge Neuron-localized Unlearning
-After selecting unforgotten samples, we localize and update only the knowledge neurons corresponding to those selected samples in the language model.
+Therefore, in each epoch’s unlearning process, we select and unlearn only questions that satisfy the knowledge memorization criteria (Described in Section 3.1).
+4.3 Knowledge-localized Unlearning
+After selecting unforgotten samples, we localize and update only the knowledge neurons corresponding to those selected samples in an LLM.
 Specifically, we first compute gradients of parameters for the selected unforgotten samples.
@@
 We finally mask gradients of the parameters for knowledge-irrelevant neurons to exclude them from the unlearning process.
-Suppose that a weight matrix $\mathbf{W}\in\mathbb{R}^{d\times k}$ is a linear matrix multiplication parameter of a language model, and the gradient computed for the parameter is $\nabla_{\mathbf{W}}\mathcal{L}=\partial\mathcal{L}/\partial\mathbf{W}$.
-Then, the gradient of $i$-th neuron (i.e., column) of the weight matrix after masking is denoted as $\nabla_{\mathbf{W}_{:,i}}\tilde{\mathcal{L}}=$ $\gamma$ $\odot$ $\nabla_{\mathbf{W}_{:,i}}\mathcal{L}$, where $\gamma\in\{\mathbf{0}_{d},\mathbf{1}_{d}\}$ and $\odot$ means the Hadamard product.
-We also can mask bias terms similar to the weight matrix.
-Notice that this method is model-agnostic since all neural networks consist of linear transformation layers.
 5 Experiments
@@
 Models.
-We adopt the instruction-tuned Gemma-2 (Gemma et al., 2024) models (2B & 9B) and the Llama-3.2 (Dubey et al., 2024) model (3B) to evaluate unlearning methods since they are among the latest open-source language models showing excellent performance.
+We adopt the instruction-tuned Gemma-2 (Gemma et al., 2024) models (2B & 9B) and the Llama-3.2 (Dubey et al., 2024) model (3B) to evaluate unlearning methods.
+These models serve as excellent starting points for unlearning evaluation, given their high default accuracy (above 80%) on the real-world entity QA benchmark.
 Data.
-We sample 5% as the forget set and 10% as the retaining set from the Base QA dataset $\mathcal{C}$ since there are generally fewer samples to unlearn than to retain in real-world scenarios.
-More experiments on varying numbers of samples for the forget set are shown in Appendix B.5.
-We select 70% of $\mathcal{C}$ as the test set, guaranteeing it is completely separate from $\mathcal{C}_{f}$ and $\mathcal{C}_{r}$.
-For the MCQA evaluation (Section 3.4), we manually select the instruction and randomly sample two false answer options from possible answers for each relation $r$.
+We sample 5% as the forget set and 10% as the retain set from the Base QA dataset since there are generally fewer samples to unlearn than to retain in real-world scenarios.
+We select 70% of $\mathcal{C}$ as the test set, guaranteeing it is completely separate from the forget and retain sets.
+For the MCQA evaluation (Section 3.4), we manually select the instruction and randomly sample two false answer options from the possible answers for each relation $r$.
 The details of an example of the MCQA format and selecting false answer options are shown in Appendix B.1 and B.2, respectively.
-We also conduct experiments on various prompt templates, and the results are described in Appendix B.6.4.
-Training Setups.
-When unlearning is applied to a language model, there is often a trade-off between unlearning knowledge (i.e., UA, UA, and MAf) and retaining the model’s overall knowledge (i.e., TA, SA, and MAt).
-Therefore, choosing the optimal model in the unlearning process is challenging since unlearning and retention are both important.
-For a fair comparison, we early stop the training procedure when UA0.33absent0.33$\leq$ 0.33≤ 0.33 is satisfied (random sampling from three options) to select the optimal model.
-More detailed experimental settings can be found in Appendix B.3.
-5.2 Baselines
-We adopt widely-used unlearning methods to evaluate the superficial unlearning: Gradient Ascent (GA), Gradient Ascent with a Retaining Loss (GAret), two Direct Preference Optimization variants (DPOmis and DPOrej), NPO (Zhang et al., 2024), and RMU (Li et al., 2024a).
-More details for the baselines are described in Appendix B.3.
-For KLUE, we select only 5% of neurons from Feed-forward networks for the knowledge neuron localization, and update them using general gradient ascent with retention loss.
+We also conduct experiments on various prompt templates, described in Appendix C.4.
+[…]
+Baselines.
+We adopt widely-used unlearning methods to assess the superficial unlearning: Gradient Ascent (GA), Gradient Ascent with a Retention Loss (GAret), two Direct Preference Optimization variants (DPOmis and DPOrej), NPO (Zhang et al., 2024), and RMU (Li et al., 2024).
+Appendix B.3 describes more details for the baselines.
+For KLUE, we select only 5% of neurons from Feed-forward networks for the knowledge localization, and update them using general gradient ascent with retention loss.
 We also use $\alpha=10$ and $N=5$ for the Superficial Knowledge Regularization term.
-The experiments analyzing various hyper-parameters are shown in Section 5.5 and Appendix B.6.
-Method UA ($\downarrow$) TA ($\uparrow$) SA ($\uparrow$) MA ($\uparrow$) Score ($\uparrow$)
+The experiments analyzing various hyper-parameters are shown in Section 5.4 and Appendix C.
+Model Method UA ($\downarrow$) TA ($\uparrow$) SA ($\uparrow$) MA ($\uparrow$) Score ($\uparrow$)
+Gemma-2
 Default 81.82 85.99 79.63 48.67 -
@@
 KLUE 36.70 82.97 74.69 58.16 69.78
-Table 2: Gemma-2 (2B) experimental results.
-We report the results of four metrics after unlearning the forget set (5%) in our settings.
+Table 2: Gemma-2 (2B) results.
+We report the results after unlearning the forget set (5%) in our settings.
 Bolded results indicate the best performance.
-We compute the accuracy over three trials and report the average accuracy.
-  Case Method Questions for Forgetting Questions for Testing Label Logit Shift
+We report the average accuracy over three trials.
+Model Method UA ($\downarrow$) TA ($\uparrow$) SA ($\uparrow$) MA ($\uparrow$) Score ($\uparrow$)
+Llama-3.2
+Default 90.91 87.28 85.65 50.57 -
+GA 35.35 54.52 39.19 52.45 52.70
+GAret 48.14 68.24 57.71 53.94 57.94
+DPOrej 46.80 69.68 55.86 54.02 58.19
+DPOmis 36.02 64.87 43.21 51.56 55.91
+KLUE 45.79 77.58 65.12 53.99 62.73
+Gemma-2
+Default 91.92 89.87 86.57 48.07 -
+GA 29.29 40.52 30.56 50.46 48.06
+GAret 45.45 83.84 68.52 50.72 64.40
+DPOrej 41.41 75.32 59.72 47.02 60.16
+DPOmis 36.36 63.15 43.06 55.45 56.32
+KLUE 40.40 89.83 81.48 60.48 72.85
+Table 3: Llama-3.2 (3B) and Gemma-2 (9B) results. We report the results unlearning the forget set (5%).
+Forget % Method UA ($\downarrow$) TA ($\uparrow$) SA ($\uparrow$) MA ($\uparrow$) Score ($\uparrow$)
+1%
+Default 72.22 85.34 71.43 54.18 -
+GA 44.44 77.80 57.14 49.43 59.98
+GAret 34.33 85.78 59.52 58.38 67.33
+DPOrej 44.44 72.84 54.76 51.79 58.73
+KLUE 36.11 85.34 63.09 59.77 68.02
+10%
+Default 83.84 85.34 76.82 50.05 -
+GA 38.38 28.02 31.13 50.41 42.79
+GAret 40.40 62.50 65.12 54.21 60.35
+DPOrej 34.85 45.26 42.38 51.29 51.02
+KLUE 40.91 81.03 69.98 59.18 67.32
+Table 4: Gemma-2 (2B) results for varying forget sample sizes (i.e., 1% and 10%). The results for 5% is also shown in Table 2
+  Case Method Questions for Forgetting Questions for Testing Label Prob Shift
 1 GAret "Where was Michael Jordan born?" (Paraphrased QA) "What city is known as the birthplace of Michael Jordan?" Brooklyn 0.5699 $\rightarrow$ 0.3333
 KLUE 0.5699 $\rightarrow$ 0.3333
-2 GAret "What is the country of citizenship of Ellen DeGeneres?" (Multi-hop QA) "What currency is associated with the country of citizenship United States dollar 0.5756 $\rightarrow$ 0.5757
+2 GAret "What is the country of citizenship of Ellen DeGeneres?" (Multi-hop QA)
+"What currency is associated with the country of citizenship
+United States dollar 0.5756 $\rightarrow$ 0.5757
 KLUE            of Ellen DeGeneres?" 0.5756 $\rightarrow$ 0.2163
@@
 Table 5: Qualitative Analysis.
-GAret and KLUE are given the same questions to forget ($\mathcal{C}_{f}$) and test ($\mathcal{C}_{p}$, $\mathcal{C}_{m}$, and $\mathcal{C}_{s}$).
+GAret and KLUE are given the same questions for forgetting and testing.
 Red texts indicate questions that should be forgotten, while blue texts should be retained.
-The "Label" and "Logit Shift" columns represent the golden labels for test questions and the logit changes corresponding to the labels, respectively.
-The underlined logit values indicate that the unlearning result is successful.
+The "Label" and "Prob Shift" columns represent the golden labels for test questions and the probability changes of the labels, respectively.
+and  indicate successful and failed unlearning, respectively.
 5.2 KLUE Mitigates Superficial Unlearning
-We investigate superficial unlearning on all baselines with Gemma-2 (2B) in the FaithUn setting, as shown in Table 2.
-First, the default Gemma-2 model can correctly answer most questions, validating that FaithUn is well constructed.
-After the unlearning process, all baselines reach UA0.33absent0.33$\leq$ 0.33≤ 0.33, which validates that all methods can unlearn target knowledge.
-However, they fail to reliably remove implicit and interconnected knowledge, suggesting that their unlearning process is superficial.
-However, our method mitigates superficial unlearning and achieves faithful unlearning compared to other baselines, without significantly damaging the other knowledge to maintain (i.e., TA, SA, and MA).
-These results demonstrate that our method accurately identifies neurons relevant to contextual knowledge and successfully erases this knowledge.
-In addition, experiments on Gemma-2 (9B) and Llama-3.2 (3B) reveal that our method outperforms baselines, with results presented in Appendix B.4.
-We also conduct ablation studies for KLUE and demonstrate the validity of our proposed methods, as detailed in Appendix B.7.
+Main experiments.
+[…]
+Forget ratios experiments.
+We conduct experiments on Gemma-2 (2B) for the varying sizes (i.e., 1%, 5%, and 10%) of the forget set to analyze the effect of unlearning samples as shown in Table 4.
+The experiments reveal that existing methods encounter more problems in unlearning when the number of forgetting samples increases, since it requires modifying a greater amount of knowledge.
+However, our proposed method consistently outperforms other baselines; thus, the performance gap between our method and the baselines widens as the number of forget samples increases.
 5.3 KLUE is Robust to Unlearning Trade-off.
@@
 However, the baselines struggle to maintain other knowledge (TA and SA) and to forget interconnected knowledge (MA).
-In contrast, KLUE demonstrates robust unlearning performance by effectively forgetting interconnected knowledge and preserving other knowledge.
+In contrast, KLUE demonstrates robust performance by effectively forgetting interconnected knowledge and preserving other knowledge.
+5.4 The Impact of Neuron Localization
+We adopt varying ratios of neuron selection $p\in\{0.01,0.05,0.1\}$ to examine the effect of the knowledge neuron on Gemma-2 (2B), shown in Figure 3.
+Also, we conduct experiments for the random neuron selection (i.e., $p\in\{0.01,0.05\}$).
+We show that KLUE achieves faithful unlearning with a neuron ratio of 0.05 or 0.1.
+In contrast, random neuron selection significantly shows superficial unlearning.
+This result reveals that the appropriate selection of knowledge neurons for unlearning is crucial to ensure the generalization of unlearning and the preservation of other knowledge.
 Figure 2: The relationship between UA and other metrics. The X-axis shows UA in descending order, and the Y-axis shows the accuracy of other metrics.
-5.5 The Impact of Neuron Localization
-We adopt varying ratios of neuron selection $p\in\{0.01,0.05,0.1\}$ to investigate the effect of the knowledge neuron on Gemma-2 (2B).
-Also, we conduct experiments for the random neuron selection (i.e., $p\in\{0.01,0.05\}$).
-As a result, we reveal that a neuron ratio of 0.05 or 0.1 contributes to achieving faithful unlearning, showing that random neuron selection more significantly triggers superficial unlearning.
-Figure 3: The ratio of neuron localization. We plot the accuracy of each metric for varying ratios of neuron localization.
+Figure 3: The ratio of neuron localization. We plot the accuracy of each metric for varying ratios of neurons.
 5.5 Qualitative Analysis
-We conduct a qualitative analysis for KLUE and GAret on Gemma-2 (2B).
-Both KLUE and GAret successfully unlearn the paraphrased question ($\mathcal{C}_{p}$), degrading label logits to 0.33.
-However, GAret has difficulty in unlearning the multi-hop question ($\mathcal{C}_{m}$), while mistakenly unlearns the same-answer questions ($\mathcal{C}_{s}$).
-On the other hand, KLUE faithfully unlearns them, mitigating superficial unlearning.
+We conduct a qualitative analysis for KLUE and GAret on Gemma-2 (2B), shown in Table 5.
+Both KLUE and GAret successfully unlearn the paraphrased question, degrading label probability to 0.33 (random guess).
+However, GAret has difficulty in multi-hop unlearning and mistakenly unlearns the same-answer questions.
+KLUE faithfully unlearns them, mitigating superficial unlearning.
+5.6 Ablation Studies
+We perform ablation studies on each KLUE method using Gemma-2 (2B) to better understand their relative importance, as shown in Table 6.
+Specifically, we remove each of the following strategies and measure the accuracy: Regularization (Section 4.1.2), Localization (Section 4.3), and Sample Selection (Section 4.2).
+The experiments demonstrate that selecting proper knowledge neurons to be updated is helpful in both handling interconnected knowledge and maintaining other knowledge.
+In addition, we reveal that Sample Selection significantly increases TA and SA, mitigating overfitting and shortcut unlearning issues.
+Module UA ($\downarrow$) TA ($\uparrow$) SA ($\uparrow$) MA ($\uparrow$) Score ($\uparrow$)
+Default 81.82 85.99 79.63 48.67 -
+KLUE 36.70 82.97 74.69 58.16 69.78
+(-) Regularization 40.40 79.74 67.59 51.24 64.54
+(-) Localization 46.46 81.68 68.52 53.51 64.31
+(-) Sample Selection 37.37 75.86 62.96 56.05 64.37
+Table 6: Ablation studies for KLUE
 6 Conclusion
-Our research identifies the limitation of existing unlearning benchmarks, which have not explored the interconnectedness of knowledge.
+Our research identifies the limitations of existing unlearning benchmarks, which have not explored the interconnectedness of knowledge.
 To overcome this issue, we define superficial unlearning and propose a new benchmark, FaithUn, for evaluating generalization, multi-hop knowledge unlearning, and shortcut unlearning.
 Using this benchmark, we empirically demonstrate that existing unlearning methods are vulnerable to superficial unlearning.
-Furthermore, we propose a novel knowledge-localized unlearning method, KLUE, and demonstrate that it outperforms existing unlearning methods, effectively mitigating superficial unlearning.
+Furthermore, we propose a novel knowledge-localized unlearning method, KLUE, which regularizes neuron attribution to identify and update only context-relevant neurons.
+We demonstrate that it outperforms existing unlearning methods, effectively mitigating superficial unlearning.
 Our paper first illuminates the phenomenon of superficial unlearning and raises a new research question for a deeper analysis of the unlearning field.
@@
 Although knowledge is more interconnected for well-known individuals, our benchmark does not examine a broader range of people.
+Our work does not evaluate knowledge editing methods, as knowledge editing and unlearning pursue different goals.
+Knowledge editing typically assesses models by post-edit accuracy, while unlearning emphasizes whether a model successfully forgets private or sensitive information.
+We therefore exclude existing knowledge editing methods from our experiments.
 Additionally, our study focuses solely on erasing the target label, leaving the issue of hallucinations in the unlearning process as future work, in line with prior studies.
@@
 Although the information of famous people is prevalent on the World Wide Web, the misuse of these data may raise ethical concerns regarding privacy.
+Acknowledgements
+This work was supported by Adobe Research.
+This work was partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [No.RS-2022-II220184, Development and Study of AI Technologies to Inexpensively Conform to Evolving Policy on Ethics & No.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University) & No.RS-2021-II212068, Artificial Intelligence Innovation Hub (Artificial Intelligence Institute, Seoul National University)].
+K.
+Jung is with ASRI, Seoul National University, Korea.
+The Institute of Engineering Research at Seoul National University provided research facilities for this work.
 Appendix A FaithUn Details
 A.1 Detailed Dataset Comparison
-In this section, we propose detailed comparisons with existing datasets to show the novelty of our benchmark clearly.
-Our benchmark aims to unlearn the knowledge of famous real-world entities, which can be prevalent in various language models, to consider the most practical situation of knowledge unlearning.
-Furthermore, our benchmark deals with the complex and interconnected nature of world knowledge; thus, we introduce three types of unlearning evaluation aspects (Generalization, Multi-hop knowledge unlearning, and Shortcut unlearning) for more deep analysis of real-world knowledge unlearning.
-In summary, MUSE, KnowUnDo, and TOFU require fine-tuning to inject knowledge before unlearning, which may reduce their practicality.
-Additionally, only RWKU and our benchmark address real-world entities as targets for unlearning.
-Furthermore, most existing benchmarks, except for RWKU and our benchmark, have not considered related knowledge.
-However, RWKU has not explored the interconnections between knowledge and shortcut unlearning problems, which have become increasingly significant in unlearning due to its reliance on a limited number of training instances.
-For example, RWKU includes a target text for unlearning: ’Please forget Stephen King, who is an American author, renowned as the "King of Horror"’.
-It also contains a related knowledge question: "’Who plays the character Jack Torrance in the film The Shining?’".
-While the two questions are somewhat related, they represent independent pieces of knowledge, as they are not interconnected like multi-hop questions.
-In conclusion, the main contribution of our benchmark lies in evaluating whether unlearning methods perform faithful unlearning while considering knowledge interconnectedness within the real-world entity unlearning setting.
-A.2 Dataset Format
-Our FaithUn benchmark includes four types of datasets: the Base QA dataset ($\mathcal{C}$), the Paraphrased QA dataset ($\mathcal{C}_{p}$), the Multi-hop QA dataset ($\mathcal{C}_{m}$), and the Same-answer QA dataset ($\mathcal{C}_{s}$).
+In this section, we present detailed comparisons with existing datasets to highlight the novelty of our benchmark.
+Our benchmark targets the unlearning of knowledge about well-known real-world entities, which are often memorized by language models, thereby addressing practical challenges in knowledge unlearning.
+Additionally, our benchmark captures the complex and interconnected nature of world knowledge by introducing three evaluation perspectives—generalization, multi-hop knowledge unlearning, and shortcut unlearning—for a more comprehensive analysis.
+In summary, MUSE, KnowUnDo, and TOFU require fine-tuning to inject knowledge prior to unlearning, which limits their practicality.
+Furthermore, existing datasets—excluding RWKU and ours—fail to evaluate whether related knowledge to the target is appropriately preserved or removed during unlearning.
+However, RWKU also has limitations in that it only evaluates knowledge through varying textual expressions (e.g., cloze test and question answering) and related but semantically disjoint facts, thus overlooking deeper relational structures.
+For example, RWKU includes a target sentence for unlearning: “Please forget Stephen King, who is an American author, renowned as the ‘King of Horror’”.
+It also presents a related question: “Who plays the character Jack Torrance in the film The Shining?”.
+RWKU evaluates whether an unlearned model preserves knowledge that is related but should not be removed after unlearning the target.
+In contrast, our benchmark introduces a more challenging setting by disentangling multiple pieces of knowledge about a single entity and evaluating whether the remaining knowledge is faithfully retained.
+We also assess the preservation of knowledge about other entities.
+These aspects are effectively evaluated by the TA and SA metrics.
+Moreover, RWKU has addressed only isolated facts with no direct knowledge dependency.
+By contrast, FaithUn is designed to evaluate unlearning in more realistic scenarios by incorporating multi-hop questions and handling directly connected pieces of knowledge.
+A.2 Details in Dataset Construction
+A.2.1 Selected Entities and Relations.
+We select 200 famous human entities and 19 relations appropriate for constructing knowledge triples from Wikidata.
+Specifically, we manually select mother, country, religion, founded by, highest point, country of citizenship, place of birth, position played on team / speciality, headquarters location, country of origin, native language, field of work, father, occupation, sport, capital, currency, location, continent as relations, which are widely-used relations to describe knowledge of human entities or other entities related to human (e.g., United States of America).
+A.2.2 Dataset Analysis.
+Dataset Format.
+Our FaithUn benchmark includes four types of datasets: the Base QA dataset, the Paraphrased QA dataset, the Multi-hop QA dataset, and the Same-answer QA dataset.
 Each instance in the Base QA dataset is matched with instances in other datasets (i.e., Paraphrased QA, Multi-hop QA, and Same-answer QA) to examine the impact of unlearning on these datasets.
@@
 Table 7: Dataset statistics.
-FaithUn includes questions to be forgotten, collectively referred to as the Base QA dataset.
-Each question in this dataset forms a cluster, and questions from other datasets (i.e., Paraphrased QA, Multi-hop QA, and Same-answer QA) are matched with those in the Base QA dataset, thereby being assigned to the corresponding cluster.
-A.3 Details in Dataset Construction
-A.3.1 Selected Entities and Relations.
-We select 200 famous human entities and 19 relations appropriate for constructing knowledge triples from Wikidata.
-Specifically, we manually select mother, country, religion, founded by, highest point, country of citizenship, place of birth, position played on team / speciality, headquarters location, country of origin, native language, field of work, father, occupation, sport, capital, currency, location, continent as relations, which are widely-used relations to describe knowledge of human entities or other entities related to human (e.g., United States of America).
-A.3.2 Dataset Analysis.
+Each question in the Base QA dataset forms a cluster, and questions from other datasets (i.e., Paraphrased QA, Multi-hop QA, and Same-answer QA) are mapped to those in the Base QA dataset, thereby being assigned to the corresponding cluster for evaluation.
 The Number of Data Instances for Each Entity.
@@
 Figure 5: Relation frequency for each dataset. the Base QA dataset (left), the Multi-hop QA dataset (middle), and the Same-answer QA dataset (right).
-A.3.3 Question Generation Prompt Templates
+A.2.3 Question Generation Prompt Templates
 We utilize GPT-4o mini to generate questions from constructed Wikidata triples, similar to (Zhong et al., 2023; Mallen et al., 2022).
@@
 Training Setups.
-We train and evaluate KLUE and other baselines on NVIDIA A100 GPU.
-For a fair comparison, we early stop the training procedure when UA0.33absent0.33$\leq$ 0.33≤ 0.33 is satisfied (random sampling from three answer options) to select the optimal model.
-Since a language model forgets all the knowledge when a learning rate is set too high, we have searched for the lowest learning rates, which can reach UA0.33absent0.33$\leq$ 0.33≤ 0.33 within the range $\lambda\in$ [1e-07, 3e-03].
-We adopt batch size $\beta=4$ for all unlearning methods.
-We compute the final loss by weighted-summing the loss of forget samples and retaining samples.
-Specifically, we use $0.7$ and $1.0$ for the loss of forget samples and the retaining samples, respectively.
-We select $e=150$ as the maximum number of epochs in the training process.
+[…]
 Baselines.
@@
 We search for $\beta_{NPO}\in[0.1,0.5]$ to optimize models.
-(5) RMU: We implement RMU (Li et al., 2024a), the representation learning-based unlearning model.
+(5) RMU: We implement RMU (Li et al., 2024), the representation learning-based unlearning model.
 For RMU experiments, we search for $\alpha_{RMU}\in\{20,50,100,150,200,300\}$ and use hyper-parameters $c=20$ and $l=7$, following the implementation details on the original GitHub Page [^3] https://github.com/centerforaisafety/wmdp.
@@
 We also use $\alpha=10$ and $N=5$ for the Superficial Knowledge Regularization term.
-The experiments analyzing varying hyper-parameters are shown in Section 5.5, Appendix B.6.2, and Appendix B.6.3.
-B.4 KLUE successfully Mitigates Superficial Unlearning on Various Language Models
-We conduct experiments on Llama-3.2 (3B) and Gemma-2 (9B) to reveal that our method is model-agnostic and generalizable.
-Table 5 shows the experimental results for Llama-3.2 (3B) and Gemma-2 (9B), respectively.
-In the Llama-3.2 (3B) results, all baselines are significantly exposed to superficial unlearning, while our method successfully outperforms other baselines in most metrics.
-Likewise, in the Gemma-2 (9B) results, the enhancement of mitigating superficial unlearning is dramatic after applying knowledge localization.
-These results demonstrate that our method can be applied to various language models and successfully mitigates superficial unlearning.
-Model Method UA ($\downarrow$) TA ($\uparrow$) SA ($\uparrow$) MA ($\uparrow$) Score ($\uparrow$)
-Llama-3.2 (3B) Default 90.91 87.28 85.65 50.57 -
-GA 35.35 54.52 39.19 52.45 52.70
-GAret 48.14 68.24 57.71 53.94 57.94
-DPOrej 46.80 69.68 55.86 54.02 58.19
-DPOmis 36.02 64.87 43.21 51.56 55.91
-KLUE 45.79 77.58 65.12 53.99 62.73
-Gemma-2 (9B) Default 91.92 89.87 86.57 48.07 -
-GA 29.29 40.52 30.56 50.46 48.06
-GAret 45.45 83.84 68.52 50.72 64.40
-DPOrej 41.41 75.32 59.72 47.02 60.16
-DPOmis 36.36 63.15 43.06 55.45 56.32
-KLUE 40.40 89.83 81.48 60.48 72.85
-Table 5: Llama-3.2 (3B) and Gemma-2 (9B) experimental results. We report the results of four metrics after unlearning the forget set (5%) in our settings. Bolded results indicate the best performance.
-B.5 KLUE is Robust to Various Forget Sample Sizes
-We conduct experiments on Gemma-2 (2B) for the varying sizes (i.e., 1%, 5%, and 10%) of the forget set to analyze the effect of unlearning samples.
-The experimental results are shown in Table 2 (5%) and Table 6 (1% and 10%).
-Our experiments reveal that existing methods undergo more problems in unlearning when the number of forget samples increases.
-Increasing the number of samples to be forgotten is more challenging since it requires modifying a greater amount of knowledge.
-However, our proposed method consistently outperforms other baselines; thus, the performance gap between our method and the baselines widens as the number of forget samples increases.
-Forget % Method UA ($\downarrow$) TA ($\uparrow$) SA ($\uparrow$) MA ($\uparrow$) Score ($\uparrow$)
-1% Default 72.22 85.34 71.43 54.18 -
-GA 44.44 77.80 57.14 49.43 59.98
-GAret 34.33 85.78 59.52 58.38 67.33
-DPOrej 44.44 72.84 54.76 51.79 58.73
-KLUE 36.11 85.34 63.09 59.77 68.02
-5% Default 81.82 85.99 79.63 48.67 -
-GA 36.02 48.92 37.19 48.34 49.61
-GAret 34.01 77.58 66.51 53.21 65.82
-DPOrej 41.75 68.96 63.58 49.67 60.11
-KLUE 36.70 82.97 74.69 58.16 69.78
-10% Default 83.84 85.34 76.82 50.05 -
-GA 38.38 28.02 31.13 50.41 42.79
-GAret 40.40 62.50 65.12 54.21 60.35
-DPOrej 34.85 45.26 42.38 51.29 51.02
-KLUE 40.91 81.03 69.98 59.18 67.32
-Table 6: Unlearning experiments for varying forget sample sizes. We report the unlearning results for the varying number of forget set (i.g., 1% and 10%). The results for 5% are also found in Table 2.
-B.6 Hyper-parameter Experiments
-B.6.1 Sequential vs. Batch Unlearning
+The experiments analyzing varying hyper-parameters are shown in Section 5.4, Appendix C.2, and Appendix C.3.
+Appendix C Additional Experiments
+C.1 Sequential vs. Batch Unlearning
 We conduct experiments on Gemma-2 (2B) to show the performance variation for varying numbers of samples unlearned in each batch.
@@
 Figure 9: The batch size experiments.
-B.6.2 Hyper-parameter ($\alpha$) Experiments
+C.2 Hyper-parameter ($\alpha$) Experiments
 We conduct hyper-parameter experiments on Gemma-2 (2B) for $\alpha\in$ {0.5, 1.0, 10.0, 20.0}, which is used to determine the magnitude of the superficial knowledge regularization, shown in Figure 10.
@@
 Figure 10: The hyper-param ($\alpha$) experiments.
-B.6.3 Neuron Ratio ($p$) Experiments
+C.3 Neuron Ratio ($p$) Experiments
 We conduct experiments on various neuron ratios to investigate the KLUE method further for Gemma-2 (2B), as shown in Table 8.
@@
 Table 8: The experiments on various neuron ratios.
-B.6.4 The Various Prompt Templates Experiments
+C.4 Various Prompt Templates Experiments
 We conduct experiments on various prompt templates to investigate the unlearning abilities of the KLUE method further for Gemma-2 (2B), as shown in Table 9.
@@
 Table 9: The experiments on different prompts.
-B.7 Ablation Studies
-We perform ablation experiments on each KLUE method using Gemma-2 2B to better understand their relative importance, as shown in Table 9.
-Regularization means the strategy of using the auxiliary regularization term for quantifying the knowledge relevance of each neuron, mitigating superficial unlearning.
-Localization corresponds to the entire knowledge neuron localization strategy.
-Sample Selection is the strategy that selects unforgotten samples by evaluating the memorization of each sample.
-For the ablation study, we remove each of them and measure the accuracy.
-As a result, we reveal that three methods significantly affect the faithfulness of unlearning.
-Regularization and Localization are useful to enhance SA and MA, mitigating superficial unlearning related to interconnected knowledge.
-These results demonstrate that selecting proper knowledge neurons to be updated is useful for handling interconnected knowledge.
-In addition, we illuminate that Sample Selection significantly increases TA and SA, mitigating overfitting and shortcut unlearning issues.
-Module UA ($\downarrow$) TA ($\uparrow$) SA ($\uparrow$) MA ($\uparrow$) Score ($\uparrow$)
+C.5 3-hop Questions Experiments
+We conduct experiments on 3-hop questions to evaluate whether unlearning methods can erase the knowledge of 3-hop questions.
+Consequently, we reveal that every method does not effectively unlearn the 3-hop knowledge.
+It is because 3-hop questions inherently form an unnatural format which is not used in practical scenarios, such as: "What is the highest point on the continent where Barack Obama holds citizenship?".
+[…]
+Default 48.67 49.70
+GA 48.34 48.32
+GAret 53.21 49.09
+RMU 53.05 47.49
+KLUE 58.16 49.66
+Table 10: Extended multi-hop (two and three hops) results for Gemma-2 (2B).
+C.6 Extended Studies on Mismatched Pairs
+We use mismatched pairs to adopt superficial knowledge regularization.
+For a deeper analysis of the regularization term, we conduct experiments on “When constructing mismatched QA pairs, why not use the same question but a randomly sampled answer?” as presented in the table below.
+We denote the new implementation using mismatched pairs of the same question but a random answer as KLUE (new).
+We observed that subtracting attribution scores based on the same question but a randomly sampled answer leads to degraded performance.
+This result suggests that when the original question is used in the knowledge regularization term, it inadvertently removes contextual knowledge rather than superficial knowledge, since it includes much of the contextual information.
+Method UA ($\downarrow$) TA ($\uparrow$) SA ($\uparrow$) MA ($\uparrow$) Score ($\uparrow$)
 Default 81.82 85.99 79.63 48.67 -
-KLUE 36.70 82.97 74.69 58.16 69.78
-(-) Regularization 40.40 79.74 67.59 51.24 64.54
-(-) Localization 46.46 81.68 68.52 53.51 64.31
-(-) Sample Selection 37.37 75.86 62.96 56.05 64.37
-Table 9: Ablation studies
+GAret 34.01 77.58 66.51 53.21 65.82
+KLUE (ours) 36.70 82.97 74.69 58.16 69.78
+KLUE (new) 45.45 81.25 70.37 52.50 62.39
+Table 11: Extended Gemma-2 (2B) studies on mismatched pairs.
+KLUE (ours) is our original model and KLUE (new) is newly introduced model by adopting $(q,a^{\prime})$ to compute the knowledge regularization.
 Type Notation Example